### v1

# FeDaL: Federated Dataset Learning for Time Series Foundation Models

Abstract
Dataset-wise heterogeneity introduces significant domain biases that fundamentally degrade generalization on Time Series Foundation Models (TSFMs), yet this challenge remains underexplored. This paper rethink the development of TSFMs using the paradigm of federated learning. We propose a novel Federated Dataset Learning (FeDaL) approach to tackle heterogeneous time series by learning dataset-agnostic temporal representations. Specifically, the distributed architecture of federated learning is a nature solution to decompose heterogeneous TS datasets into shared generalized knowledge and preserved personalized knowledge. Moreover, based on the TSFM architecture, FeDaL explicitly mitigates both local and global biases by adding two complementary mechanisms: Domain Bias Elimination (DBE) and Global Bias Elimination (GBE). FeDaL’s cross-dataset generalization has been extensively evaluated in real-world datasets spanning eight tasks, including both representation learning and downstream time series analysis, against 54 baselines. We further analyze federated scaling behavior, showing how data volume, client count, and join rate affect model performance under decentralization.
1 Introduction
Time series data underpin critical decisions across domains such as climate bi2023accurate , traffic deng2024multi , and healthcare wang2024medformer , where tasks like forecasting lim2021time , imputation chen2024federated , classification zhang2025timesbertbertstylefoundationmodel , and anomaly detection liu2024timer are essential for operational intelligence. While early work emphasized task-specific models tailored to narrow objectives zhou2022fedformer , recent advances have shifted toward general-purpose time series foundation models (TSFMs) liu2024timer ; das2023decoder , inspired by the success of large language models. These TSFMs aim to generalize across various tasks and domains, but doing so requires access to both diverse public datasets and large-scale private data. Federated Learning (FL) mcmahan2017communication presents a scalable framework for training TSFMs collaboratively chen2024federated , as it enables training across distributed data while preserving privacy and reducing the computational burden associated with centralized training ye2024openfedllm .
Figure 1: Key domain biases in heterogeneous time series: temporal resolution bias, physical constraint bias, and pattern transition bias. Information Lost: Low-resolution sequences encode less detail within a fixed lookback window. Pattern Break: Changes at specific time points (or segments). FL Settings: FedAvg-Independent (each dataset as a client), FedAvg-Mixed (datasets mixed and split into two clients), Stand-alone (no aggregation). Other settings follow Section 4. Lower MSE indicates better performance. Best viewed in color and zoomed in.
Despite its promise, training TSFMs via FL introduces unique challenges, as robust generalization requires addressing domain biases caused by heterogeneous datasets, with biases that are both diverse and underexplored in prior work chen2024federated . Unlike text/images, time series vary widely in resolution, physical constraints, and temporal dynamics, complicating the extraction of domain-invariant patterns. Although FL inherently supports heterogeneity through its decentralized nature, existing works chen2024federated primarily focus on aligning trends across domains, overlooking deeper, dataset-specific structural biases that can severely hinder generalization. We identify three key types of such bias (Figure˜1): (1) Temporal resolution bias, sequences with different sampling rates encode inconsistent contextual information under a fixed window (e.g., 120 steps cover five days in hourly weather data but only two hours in minute-level energy data); (2) Physical constraint bias, domain-specific dynamics governed by unrelated physical laws (e.g., temperature variation vs. electric current) reduce representational transferability; and (3) Pattern transition bias, initially similar trends across domains may diverge sharply due to exogenous events (e.g., traffic vs. web visits), breaking assumptions of shared temporal structures. These biases pose fundamental obstacles to learning robust and cross-domain transferable patterns via FL, ultimately limiting the generalization capacity of the learned TSFMs.
This paper introduces Federated Dataset Learning (FeDaL), a brand-new framework for training Time Series Foundation Models under federated settings, with the goal of capturing generalized temporal patterns from heterogeneous datasets. To address domain-level heterogeneity, FeDaL incorporates two complementary components: Domain Bias Elimination (DBE) and Global Bias Elimination (GBE), operating from the client and server perspectives, respectively. DBE approximates and disentangles client-specific biases in a context-agnostic manner by explicitly modeling and storing local bias representations. GBE enhances global model generalization by aligning client contributions through gradient-level correction and privacy-preserving, fine-grained server-side tuning, enabling adaptive bias mitigation during aggregation. Together, these components allow the pretrained TSFM to learn domain-invariant temporal representations, facilitating strong cross-domain generalization while maintaining privacy and scalability. Our key contributions are summarized as follows:
  • We identify three key domain biases (temporal resolution, physical constrains, and pattern transition) that hinder the learning of general temporal patterns in FL-based TSFM training.
  • We propose Federated Dataset Learning (FeDaL), a novel framework for learning TSFMs in federated settings, which explicitly addresses dataset heterogeneity through context-agnostic Domain Bias Elimination (DBE) and server-side Global Bias Elimination (GBE).
  • We perform the first systematic analysis of TSFM scaling behaviors from a federated learning perspective, offering empirical insights into how data volume, client population, and participation rate affect model generalization under decentralized settings.
  • Extensive experiments on real-world time series demonstrate that FeDaL enables strong in-domain and out-of-distribution generalization, consistently outperforming state-of-the-art baselines across forecasting, imputation, classification, and anomaly detection tasks.
2 Preliminaries
Pre-trained Time Series Models
Pretraining has emerged as a key strategy for building effective time series models, supporting both task-specific nie2022time ; jin2023time and general-purpose TSFMs goswami2024moment ; woo2024unifiedtraininguniversaltime ; shi2024time . Most existing TSFMs adopt masked reconstruction at the time-point zeng2023transformers or patch level liu2024timer ; garza2023timegpt , achieving strong zero-shot performance. Recent efforts such as Moment goswami2024moment , Moirai woo2024unifiedtraininguniversaltime , TimeFM das2024decoder , and Chronos ansari2024chronos push this further with large-scale masked pretraining to enhance downstream adaptation. However, these models are often tailored to forecasting and show limited generalization to tasks like classification or anomaly detection. Moreover, their reliance on centralized, large-scale datasets restricts real-world applicability, where data is often siloed and heterogeneous chen2024federated ; zhuang2023foundation . Federated learning mcmahan2017communication offers a practical alternative by enabling collaborative model training across decentralized domains while preserving privacy chen2024federated . Building on this paradigm, we propose a brand-new federated framework for robust TSFM pretraining under decentralized and domain-heterogeneous settings, enabling broad adaptation across downstream tasks and domains.
Federated Learning in Time Series
FL provides a practical framework for collaborative training without centralized data fusion mcmahan2017communication , especially important given rising privacy concerns and the emergence of foundation models zhuang2023foundation . In time series analysis, FL has enabled both personalized models liu2024time and, more recently, generalized TSFMs chen2024federated , drawing increasing attention chen2023prompt ; chen2023federated ; sun2024hifi . While existing methods focus on client-level heterogeneity, they often overlook domain-level bias—a key challenge in time series where clients frequently hold non-overlapping, domain-specific data. These domain biases result in inconsistent optimization paths and poorly aligned latent patterns, limiting TSFM generalization. To address this, we introduce FeDaL, a brand-new FL paradigm that explicitly models and aligns domain-specific and global representations. By mitigating domain bias, FeDaL enables the training of TSFMs with improved cross-domain generalization.
Problem Definition
We consider a FL setting with a server and $N$ clients, where each client $i$ owns a local time series dataset ${\mathcal{D}}_{i}$ drawn from one or more domains. These datasets are inherently heterogeneous due to variations in sampling rates, domain-specific characteristics, and temporal dynamics. The objective is to collaboratively train a unified TSFM that generalizes across clients and supports diverse downstream tasks. Formally, we aim to minimize the global objective: $F(\theta):=\operatorname*{arg\,min}_{\theta}\sum_{i=1}^{N}\frac{n_{i}}{n}F_{i}(\theta_{i};{\mathcal{D}}_{i})$, where $F_{i}(\cdot)$ denotes the local loss, and $_{i}, nn$ are the local and total sample counts, respectively. This work asks: Can we train a TSFM under federated constraints that captures cross-domain temporal structure while minimizing the impact of domain-specific biases?
3 Methodology: Federated Dataset Learning
Overview
Heterogeneous time series datasets often introduce domain biases, causing local models to overfit to spurious patterns tied to particular datasets or clients. When these locally biased models are aggregated, their misaligned representations and divergent optimization trajectories result in a globally biased TSFM with limited generalization. These challenges are especially pronounced in massive, high-dimensional, and structurally diverse time series data. To address this, we propose two complementary mechanisms—Domain Bias Elimination (DBE) and Global Bias Elimination (GBE)—that operate at the client and server levels, respectively. Together, they explicitly mitigate domain-level bias and enable the alignment of complex temporal patterns into a unified, generalizable representation. The detailed algorithmic implementation of our FeDaL (Left, Figure˜2) is provided in Algorithm˜1. The remainder of this section details the designs and roles of DBE and GBE.
Figure 2: (Left) FedDaL with two clients example: DBE alleviates domain bias from the client perspective, while GBE addresses global bias from the server side. (Right) Each client employs a Decoder-only Transformer liu2024time ; liu2024timer , extracting trend and season biases from latent representations via plug-and-play DBE block.
3.1 Domain Bias Elimination
Motivation
When client models overfit to domain-specific artifacts, their latent representations tend to entangle general temporal features with spurious local biases, undermining the quality of representation learning and hindering downstream generalization. To mitigate this, we introduce Domain Bias Elimination (DBE) (refer to  Right, Figure˜2). DBE decomposes the latent representation of masked input patches into context-agnostic components (trend and seasonal), from which a trainable local bias vector is estimated. This vector is injected back during reconstruction to isolate and suppress domain-specific biases. Additionally, we regularize its alignment to a global bias reference, guiding the model toward disentangling spurious local patterns from generalizable temporal structure.
Given a masked input sequence ${\bm{X}}_{i}^{\text{masked}}$ on client $i$, its latent representation is obtained using the local backbone $f_{\theta^{b}_{i}}(\cdot)$. We introduce a plug-and-play module, the DBE block, which incorporates time series decomposition wu2021autoformer and averaging to explicitly estimate local bias. Specifically, we first decompose the masked input representations into trend and seasonal components, as follow:
$\mathbf{h}_{i}^{\text{trend}},\mathbf{h}_{i}^{\text{season}}=\texttt{Decomp}(f_{\theta^{b}}({\bm{X}}_{i}^{\text{masked}}),\tau),\quad\mathbf{b}_{i}^{\text{season}},\mathbf{b}_{i}^{\text{trend}}=\texttt{Mean}(\mathbf{h}_{i}^{\text{season}}),\texttt{Mean}(\mathbf{h}_{i}^{\text{trend}})$ (1)
where $\tau$ denotes the decomposition granularity. The components $\mathbf{h}_{i}^{\text{trend}},\mathbf{h}_{i}^{\text{season}}$ capture generalizable, context-independent temporal patterns. By averaging these components, we extract persistent directional deviations in the latent space zhang2023eliminating , which serve as compact and effective approximations of client-specific biases arising from the accumulation of complex time series dataset heterogeneity. We estimate the local bias vector by summing the two components: $\hat{\mathbf{b}}_{i}^{p}=\mathbf{b}^{\text{trend}}_{i}+\mathbf{b}^{\text{season}}_{i}$. Both $\mathbf{b}_{i}^{\text{trend}}$ and $\mathbf{b}_{i}^{\text{season}}$ are kept trainable to adaptively capture evolving client-specific characteristics. During training, the model reconstructs the original input by injecting the bias vector into the latent space:
${\mathcal{L}}_{i}(\theta_{i}):=\mathbb{E}\left[\left\|f_{\theta^{h}}(f_{\theta^{b}}({\bm{X}}^{\text{masked}})+\hat{\mathbf{b}}_{i}^{p})-{\bm{X}}^{\text{orig}}\right\|^{2}\right].$ (2)
Here, $\theta_{i}=[\theta^{b},\theta^{h},\hat{\mathbf{b}}_{i}^{p}]$ includes all local learnable parameters, and ${\bm{X}}_{i}^{\text{orig}}$ is the original input sequence. To promote alignment across clients and reduce divergence caused by heterogeneous biases, we introduce an explicit regularization term based on bias approximation to rewrite Eq. 2 as:
${\mathcal{L}}_{i}(\theta_{i}):=\mathbb{E}[\|f_{\theta^{h}}(f_{\theta^{b}}({\bm{X}}^{\text{masked}})+\hat{\mathbf{b}}_{i}^{p})-{\bm{X}}^{\text{orig}}\|^{2}]+\lambda\|\hat{\mathbf{b}}_{i}^{p}-\mathbf{b}^{g}\|^{2}.$ (3)
This encourages disentanglement of biased and unbiased components in the local representation, guiding the projection head to emphasize generalizable features. The global bias vector $\mathbf{b}^{g}$ is computed once during initialization (refer to Line 3 in Algorithm 1), and alignment is enforced via a mean squared error (MSE) penalty between the local and global biases. The regularization strength is controlled by a tunable coefficient $\lambda$, optimized jointly via SGD. Since the bias term in Eq. 3 depends on aggregated statistics across the full local dataset—while SGD operates on mini-batches—we approximate the global signal using an exponential moving average zhang2015deep across batches:
$\mathbf{b}_{i}^{\text{trend}}=(1-\mu)\mathbf{b}_{i,\text{old}}^{\text{trend}}+\mu\mathbf{b}_{i,\text{new}}^{\text{trend}},\quad\mathbf{b}_{i}^{\text{season}}=(1-\mu)\mathbf{b}_{i,\text{old}}^{\text{season}}+\mu\mathbf{b}_{i,\text{new}}^{\text{season}},$ (4)
where $\mu$ is a smoothing coefficient, and “new“ and “old“ denote current and previous batch-wise estimates, respectively. This allows stable bias tracking throughout local training.
3.2 Global Bias Elimination
Motivation
While DBE effectively addresses spurious correlations at the client level in cross-domain time series, it remains insufficient for correcting global biases that emerge from heterogeneous temporal patterns and domain shifts across clients. For this, we propose Global Bias Elimination (GBE), a dual-component method that operates from the server side to mitigate representation divergence and improve global alignment. GBE comprises two components: (1) Representation Dynamic Correction, which corrects client-server gradient drift through a server-maintained correction term; and (2) Server-side Core-set Tuning, which refines the global model using privacy-preserving, client-selected data summaries. Together, these components address persistent cross-client discrepancies, allowing the global model to learn domain-invariant patterns while maintaining privacy.
Representation Dynamic Correction
Inspired by acar2021federated , we maintain a server-side state vector $\mathbf{s}$ that accumulates client-server gradient differences across communication rounds. This state serves as a correction term during global aggregation, helping to stabilize convergence and counteract representation drift. The state vector generation and global correction can be expressed as:
${\color[rgb]{0,0,1}\text{[State Updating] }}\mathbf{s}^{r}=\mathbf{s}^{r-1}-\beta\sum_{i}(\theta_{i}^{r}-\theta^{g,r-1})\quad{\color[rgb]{1,0,0}\text{[Global Correction]}}\quad\theta^{\prime\prime g,r}=\theta^{\prime g,r}-\frac{1}{\beta}\mathbf{s}^{r}$ (5)
Here, the initial state vector $s^{0}$ is initialized using the global model $\theta^{g,0}$, $\beta$ controls the strength of the correction, and $\theta^{\prime g,r}$ is the initial global model at the $r$-th round, obtained by weighted averaging mcmahan2017communication . By dynamically adjusting global updates in each round, this mechanism mitigates the impact of client-specific deviations and promotes a more generalized global representation.
Server-side Core-set Tuning
To further mitigate global bias, we introduce a server-side fine-tuning stage based on compact core-sets constructed locally by clients. Each core-set aims to approximate the client’s data distribution via gradient matching zhao2020dataset , selecting a small subset whose gradients resemble those of the full dataset. Given a local dataset $\mathcal{X}=\{x_{1},\dots,x_{Z}\}$, each client randomly samples a mini-batch $\tilde{\mathcal{X}}\subset\mathcal{X}$ of size $K\ll Z$ for efficiency and privacy. The objective is to construct an initial core-set $\mathcal{C}_{\text{init}}$ by minimizing the gradient discrepancy between $\mathcal{C}_{\text{init}}$ and $\tilde{\mathcal{X}}$:
$\mathcal{L}_{\text{match}}({\mathcal{C}}_{\text{init}})=\sum\nolimits_{x_{k}\in\tilde{\mathcal{X}}}\left|\nabla_{\theta}f_{\theta}(\mathcal{C}_{\text{init}})-\nabla\theta f_{\theta}(x_{k})\right|_{2}^{2},$ (6)
where $f_{\theta}$ denotes the current model on clients. The core-set is iteratively updated via: $\mathcal{C}_{\text{init}}\leftarrow\mathcal{C}_{\text{init}}-\eta\nabla_{\mathcal{C}_{\text{init}}}\mathcal{L}_{\text{match}}$, with learning rate $\eta$. To preserve privacy, we apply a Fourier-based perturbation mechanism that masks raw sequence details while retaining semantic structure. This is motivated by the observation that the phase of the Fourier transform encodes high-level temporal semantics, whereas the amplitude captures fine-grained variations. Accordingly, we perturb only the amplitude in the frequency domain: ${\mathcal{C}}^{\prime}={\mathcal{F}}^{-1}({\mathcal{F}}(\mathcal{C}_{\text{init}})+\epsilon{\mathcal{N}}(0,1))$, where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the forward and inverse Fourier transforms, and $\epsilon$ controls the noise intensity. To ensure that $\mathcal{C}^{\prime}$ remains semantically faithful to the original data, we align its latent representation with that of the sampled mini-batch $\tilde{\mathcal{X}}$:
$\mathcal{L}_{\text{align}}({\mathcal{C}}^{\prime})=\frac{1}{K}\|\sum\nolimits_{c_{j}\in\mathcal{C}^{\prime}}f_{\theta}(c_{j})-\sum\nolimits_{x_{k}\in\tilde{\mathcal{X}}}f_{\theta}(x_{k})\|_{2}^{2}.$ (7)
The semantically aligned core-set $\mathcal{C}$, optimized from the perturbed $\mathcal{C}^{\prime}$ via Eq. 7, is then uploaded to the server. The server uses $\{\mathcal{C}_{i}\}^{N}$ to perform targeted fine-tuning on the global model, following the process of masked reconstruction nie2022time , yielding an refined version $\theta^{gt,r}$. To mitigate catastrophic forgetting and maintain continuity with prior global knowledge, we apply convex model fusion marfoq2022personalized :
$\theta^{g,r}=\alpha\theta^{\prime\prime g,r}+(1-\alpha)\theta^{gt,r},$ (8)
where $\alpha\in[0,1]$ is a fusion coefficient that controls the trade-off between stability and plasticity.
Algorithm 1 Implementation of our Federated Dataset Learning (FeDaL)
1:Clients $\{c_{i}\}_{i=1}^{N}$, rounds $R$, local epochs $E$, core-set size $K$, period $\tau$; hyperparameters $\lambda$, $\alpha$, $\epsilon$.
2:Time Series Foundation Model $\theta^{g}=\{\theta^{b},\theta^{h}\}$
3:// Initialization Phase:
4:Server broadcasts $\theta^{g,0}=\{\theta^{b,0},\theta^{h,0}\}$ to initialize all clients; sets server state $\mathbf{s}=\mathbf{0}$.
5:Initialize global bias $\mathbf{b}^{g}$ by weighted averaging client biases $\hat{\mathbf{b}}^{p}_{i}$} after one warm-up epoch: $\mathbf{b}^{g}=:\sum_{i}\frac{n_{i}}{n}\hat{\mathbf{b}}^{p}_{i}$.
6:Client $i$ sets $\mathbf{b}_{i}^{\text{trend},0}=\mathbf{0}$, $\mathbf{b}_{i}^{\text{season},0}=\mathbf{0}$, and receives $\hat{\mathbf{b}}_{i}^{p,0}$.
7:for round $r=1$ to $R$ do
8:  Server samples client subset ${\mathcal{S}}_{r}$ with ratio $\rho$, and sends $\theta^{b,r-1},\theta^{h,r-1}$} to each client in ${\mathcal{S}}_{r}$
9:  // Local Updating Phase:
10:  for client $c_{i}\in\mathcal{S}_{r}$ in parallel do
11:    Obtain $\mathbf{h}_{i}^{\text{trend}},\mathbf{h}_{i}^{\text{season}}$ using time series decomposition $\triangleright$ Refer to Eq. 1
12:    Update local bias $\hat{\mathbf{b}}_{i}^{p,r}=\texttt{Mean}(\mathbf{h}_{i}^{\text{trend}})+\texttt{Mean}(\mathbf{h}_{i}^{\text{season}})$
13:    Optimize $\{\theta_{i}^{b,r},\theta_{i}^{h,r},\hat{\mathbf{b}}^{p,r}\}$ via SGD for $\min_{\theta_{i}}{\mathcal{L}}_{{\mathcal{D}}_{i}}(\theta_{i})$ with $\lambda$ for $E$ epoch $\triangleright$ Refer to Eq. 3
14:    Obtain core-set ${\mathcal{C}}_{i}$ via SGD for $\min{\mathcal{L}}_{\text{match}}({\mathcal{C}}_{\text{init}})$ and $\min{\mathcal{L}}_{\text{align}}({\mathcal{C}}^{\prime})$ with $K,\eta$ $\triangleright$ Refer to Eqs. 67
15:    Upload $\{\theta_{i}^{b,r},\theta_{i}^{h,r},{\mathcal{C}}_{i}\}$ to the server
16:  end for
17:  // Server Aggregation Phase:
18:  Update global model $\theta^{g,r}=\sum_{i}\frac{|\mathcal{D}_{i}|}{\sum_{j}|\mathcal{D}_{j}|}\theta_{i}^{r}$; Update server state $\mathbf{s}^{r}=\mathbf{s}^{r-1}-\alpha\sum_{i}(\theta_{i}^{r}-\theta^{g,r-1})$
19:  Obtain refined global model $\theta^{\text{gt},r}$ by fine-tuning on the collection of core-sets $(\bigcup_{i}\mathcal{C}_{i})$
20:  Update the TSFM $\theta^{g,r}=\alpha\cdot\theta^{\prime\prime g,r}+(1-\alpha)\cdot\theta^{\text{gt},r}$, Update global bias $\mathbf{b}^{g,r}=\frac{1}{|\mathcal{S}_{r}|}\sum_{i\in\mathcal{S}_{r}}\mathbf{b}^{p,r}_{i}$
21:end for
4 Experiments and Results
In this section, we comprehensively evaluate FeDaL across three dimensions: federated representation learning, downstream generalization, and federated scaling behavior. Our experiments demonstrate that FeDaL (i) learns domain-agnostic representations in highly heterogeneous settings, (ii) enables strong generalization across forecasting, imputation, classification, and anomaly detection tasks, and (iii) exhibits favorable scaling properties under increasing data, client count and join ratio.
4.1 Federated Time Series Representation Learning
Setup.
Learning dataset-invariant patterns from unlabeled and heterogeneous time series is critical for advancing TSFMs in federated settings. We evaluate FeDaL on two large-scale benchmarks: UTSD liu2024timer (7 domains, 1B) with domain-mixed (DM) partitioning, and CTSD chen2024federated (6 domains, 500M) with domain-independent (DI) partitioning. In DM, each client sees data from multiple domains, whereas in DI, clients are domain-specific. We compare FeDaL with five FL baselines, including FedAvg mcmahan2017communication , FedProx li2020federated , FedPer arivazhagan2019federated , FedRep collins2021exploiting , and TSFM-specific FFTS chen2024federated . All models are trained to reconstruct masked patches in an unsupervised manner nie2022time , using input length 1024, patch length 32, and batch size 1024. We use a masking ratio of 75% during training and report reconstruction MSE across varying mask ratios $\{20\%,35\%,50\%,75\%,90\%\}$ to assess robustness. The client join ratio $\rho=70\%$, with hyperparameters set as: core-set size $K=1024$, decomposition period $\tau=4$, and $\lambda=10^{-2}$, $\alpha=0.7$, and $\beta=10^{-1}$ for DBE and GBE. Each method is trained for 200 rounds. Full dataset statistics and implementation details are provided in Appendix˜B.
Table 1: Federated representation learning results under varying patch masking ratios. For UTSD, we simulate two levels of domain heterogeneity (H1 and H2) as detailed in the Appendix˜B. For CTSD, each dataset is treated as a domain-independent client. Bold indicates the best result, Underline the second best. $\dagger$} denotes evaluation using personalized models after server averaging; $\ddagger$} indicates client-side evaluation without aggregation.
Method
USTD Dataset (10 Clients, DM)
CTSD Dataset (18 Clients, DI)
20%
35%
50%
75%
90%
20%
35%
50%
75%
90%
H1 H2 H1 H2 H1 H2 H1 H2 H1 H2
FedAvg 0.387 0.404 0.473 0.489 0.550 0.565 0.636 0.602 0.882 0.901 0.350 0.388 0.405 0.480 0.652
FedProx 0.382 0.401 0.469 0.486 0.546 0.553 0.638 0.603 0.880 0.889 0.336 0.390 0.400 0.454 0.638
FedPer 0.343 0.385 0.459 0.578 0.540 0.533 0.621 0.580 0.860 0.863 0.330 0.381 0.395 0.439 0.606
FedRep 0.390 0.413 0.451 0.560 0.529 0.537 0.621 0.592 0.853 0.860 0.352 0.373 0.386 0.440 0.598
FFTS 0.333 0.327 0.450 0.410 0.526 0.510 0.630 0.584 0.870 0.823 0.300 0.357 0.379 0.436 0.610
Stand-alone 0.376 0.380 0.454 0.450 0.535 0.521 0.643 0.616 0.875 0.870 0.342 0.381 0.395 0.470 0.646
FeDaL (Ours) 0.348 0.300 0.436 0.422 0.521 0.489 0.596 0.549 0.852 0.795 0.310 0.319 0.343 0.405 0.560
Main Results.
The main results of representation learning are presented in Table˜1, our FeDaL consistently outperforms baselines across varying heterogeneous settings. On USTD, it achieves superior results under both levels of domain heterogeneity (H1 and H2), and on CTSD, it excels across all masking ratios under domain-independent partitioning. Compared to the state-of-the-art federated TSFM pretraining method FFTS, FeDaL reduces reconstruction MSE by an average of 4.16% on USTD and 8.86% on CTSD, with particularly notable gains at higher masking ratios. These highlight our FeDaL’s ability to balance domain-specific feature learning on clients with domain-invariant aggregation at the server, leading to more generalizable representations under significant domain bias.
Figure 3: Results of ablation (Left) and hyperparameter sensitively (Right) on federated representation learning. Results on the UTSD dataset are averaged over H1 and H2. $\downarrow$ indicates performance drop; $\uparrow$, improvement relative to the original FeDaL. Experimental settings follow the main setup. Best viewed in color and with zoom.
Ablation and Hyperparameter Sensitively Results
We analyze the contribution of key components and the robustness of FeDaL to hyperparameter choices (Figure˜3). Ablations reveal that removing DBE or GBE significantly harms representation quality, confirming their role in addressing domain heterogeneity. Disabling DBE’s alignment step (Eq. 3) further degrades performance, emphasizing the importance of bias alignment. Removing Core-set or Correction (Eq. 5) impairs global adaptation, showing their necessity for model refinement under heterogeneity. Sensitivity results show: (i) overly large alignment weight $\lambda$ over-constrains local representations; (ii) larger core-sets (e.g., 2048, 3072) bring marginal gains at the cost of privacy and communication; (iii) extreme fusion weights $\alpha$ weaken generalization; (iv) unstable $\beta$ values impair client-server representation blending.
Bias Representation Analysis
Figure 4: Local bias change across communication rounds (R1 - R60) for select clients (Clients 3, 7, 9).
To evaluate how effectively our FeDaL reduces domain bias during federated training, we visualize the evolution of local bias representations across three representative clients (Client 3, 7, 9) from round 1 to round 60, as shown in Figure˜4. At the initial round (R1), substantial variation is observed both across dimensions and among clients, reflecting strong intra- and inter-client biases due to limited generalization. As training progresses (e.g., R5 and R20), these biases progressively diminish. By round 60 (R60), the representations across clients exhibit strong alignment, indicating that FeDaL gradually encourages convergence toward domain-invariant features. This provides intuitive evidence that FeDaL effectively mitigates client-specific bias and facilitates the learning of generalized, cross-domain time series representations.
4.2 Downstream Task Adaption as Time Series Foundation Models
To evaluate the generality of FeDaL-pretrained TSFMs, we test on diverse downstream tasks: long-/short-term forecasting, imputation, classification, and anomaly detection. Pretraining is conducted on LOTSA dataset woo2024unifiedtraininguniversaltime  (231B time points across 174 client-specific subsets spanning nine domains). We use 1000 rounds with 10 local steps, and SGD with learning rate $1e^{-2}$. Other hyperparameters follow Section˜4.1, and data preprocessing details are provided in Appendix˜B, full results in Appendix˜C.
Table 2: Long-term forecasting results (averaged across horizons $\{96,192,336,720\}$ for ETT-series and Weather, and $\{24,36,48,60\}$ for ILI) in full-shot and few-shot settings (5% and 10% training data). Bold: best; Underline: second best. None of these datasets were included in pretraining. * denotes a “former“ suffix.
Models FeDaL (Ours) FFTS FedAvg TimeMixer Time-LLM GPT4TS PatchTST TimesNet DLinear Fed.* Auto.* Stationary LightTS In.* Re.*
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Full-shot Long-term Forecasting
ETTh1 0.380 0.409 0.391 0.412 0.412 0.431 0.448 0.443 0.408 0.423 0.465 0.455 0.413 0.430 0.458 0.45 0.422 0.437 0.440 0.460 0.496 0.487 0.570 0.537 0.491 0.479 1.040 0.795 1.029 0.805
ETTh2 0.334 0.377 0.334 0.389 0.340 0.382 0.364 0.394 0.334 0.383 0.381 0.412 0.330 0.379 0.414 0.427 0.431 0.446 0.437 0.449 0.450 0.459 0.526 0.516 0.602 0.543 4.431 1.729 6.736 2.191
ETTm1 0.319 0.365 0.323 0.374 0.333 0.367 0.381 0.395 0.329 0.372 0.388 0.403 0.351 0.380 0.400 0.406 0.357 0.378 0.448 0.452 0.588 0.517 0.481 0.456 0.435 0.437 0.961 0.734 0.799 0.671
ETTm2 0.261 0.319 0.253 0.314 0.254 0.316 0.275 0.323 0.251 0.313 0.284 0.339 0.255 0.315 0.291 0.333 0.267 0.333 0.305 0.439 0.327 0.371 0.306 0.347 0.409 0.436 1.410 0.810 1.479 0.915
Weather 0.213 0.255 0.217 0.256 0.226 0.260 0.241 0.272 0.225 0.257 0.237 0.270 0.255 0.264 0.259 0.287 0.248 0.300 0.309 0.360 0.338 0.382 0.288 0.314 0.261 0.312 0.634 0.548 0.803 0.656
ILI 1.355 0.773 1.389 0.798 1.410 0.800 2.039 0.899 1.435 0.801 1.925 0.903 1.443 0.797 2.139 0.931 2.169 1.041 2.847 1.144 3.006 1.161 2.077 0.914 7.382 2.003 5.137 1.544 4.724 1.445
Few-shot Long-term Forecasting (5%)
ETTh1 0.592 0.511 0.601 0.520 0.630 0.537 1.193 0.740 0.627 0.543 0.681 0.560 0.694 0.569 0.925 0.647 0.750 0.611 0.658 0.562 0.722 0.598 0.943 0.646 1.451 0.903 1.225 0.817 1.241 0.835
ETTh2 0.350 0.391 0.360 0.400 0.388 0.416 0.502 0.493 0.382 0.418 0.400 0.433 0.827 0.615 0.439 0.448 0.694 0.577 0.463 0.454 0.441 0.457 0.470 0.489 3.206 1.268 3.922 1.653 3.527 1.472
ETTm1 0.400 0.401 0.417 0.415 0.415 0.414 0.566 0.503 0.425 0.434 0.472 0.450 0.526 0.476 0.717 0.561 0.400 0.417 0.730 0.592 0.796 0.620 0.857 0.598 1.123 0.765 1.163 0.791 1.264 0.826
ETTm2 0.240 0.307 0.253 0.299 0.258 0.308 0.340 0.371 0.274 0.323 0.308 0.346 0.314 0.352 0.344 0.372 0.399 0.426 0.381 0.404 0.381 0.404 0.341 0.372 1.415 0.871 3.658 1.489 3.581 1.487
Weather 0.222 0.286 0.241 0.295 0.242 0.295 0.260 0.303 0.260 0.309 0.263 0.301 0.269 0.303 0.298 0.318 0.263 0.308 0.309 0.353 0.309 0.353 0.327 0.328 0.305 0.345 0.584 0.527 0.447 0.453
Few-shot Long-term Forecasting (10%)
ETTh1 0.540 0.505 0.549 0.513 0.581 0.520 0.613 0.520 0.556 0.522 0.590 0.525 0.633 0.542 0.869 0.628 0.691 0.600 0.639 0.561 0.702 0.596 0.915 0.639 1.375 0.877 1.199 0.808 1.249 0.833
ETTh2 0.352 0.388 0.360 0.391 0.382 0.400 0.402 0.433 0.370 0.394 0.397 0.421 0.415 0.431 0.479 0.465 0.605 0.538 0.466 0.475 0.488 0.499 0.462 0.455 2.655 1.160 3.872 1.513 3.485 1.486
ETTm1 0.389 0.407 0.389 0.410 0.410 0.428 0.487 0.461 0.404 0.427 0.464 0.441 0.501 0.466 0.677 0.537 0.411 0.429 0.722 0.605 0.802 0.628 0.797 0.578 0.971 0.705 1.192 0.821 1.426 0.856
ETTm2 0.261 0.312 0.264 0.315 0.289 0.339 0.311 0.367 0.277 0.323 0.293 0.335 0.296 0.343 0.320 0.353 0.316 0.368 0.463 0.488 1.342 0.930 0.332 0.366 0.987 0.756 3.370 1.440 3.978 1.587
Weather 0.207 0.256 0.214 0.261 0.240 0.278 0.242 0.281 0.234 0.273 0.238 0.275 0.242 0.279 0.279 0.301 0.241 0.283 0.284 0.324 0.300 0.342 0.318 0.323 0.289 0.322 0.597 0.495 0.546 0.469
Average 0.396 0.390 0.405 0.396 0.422 0.404 0.578 0.435 0.419 0.406 0.478 0.427 0.484 0.434 0.577 0.469 0.534 0.469 0.607 0.505 0.708 0.513 0.620 0.491 1.562 0.766 2.142 1.010 2.359 1.059
$1^{st}$ Count 30 2 0 0 2 0 1 0 1 0 0 0 0 0 0
Long-Term Forecasting
[…]
Table 3: Average zero-shot forecasting performance across horizons $\{96,192,336,720\}$ for observation lengths $\{512,1024,2048,3072\}$. Bold: best; Underline: second best. Full results in Table˜14, Appendix˜C.
Models Federated Learning Methods Pretrained Time Series Foundation Models
Metrics FeDaL (Ours) FFTS FedAvg Moiraismall Moiraibase Moirailarge TimesFM Moment Chronossmall Chronosbase Chronoslarge
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.407 0.429 0.425 0.437 0.438 0.449 0.428 0.427 0.417 0.419 0.480 0.439 0.473 0.443 0.683 0.566 0.545 0.472 0.591 0.468 0.588 0.466
ETTh2 0.361 0.382 0.370 0.374 0.390 0.401 0.361 0.384 0.362 0.382 0.367 0.377 0.392 0.406 0.361 0.409 0.424 0.430 0.405 0.410 0.455 0.427
ETTm1 0.360 0.390 0.364 0.420 0.378 0.410 0.436 0.410 0.406 0.385 0.422 0.391 0.433 0.418 0.670 0.536 0.640 0.499 0.645 0.500 0.555 0.465
ETTm2 0.292 0.341 0.317 0.362 0.322 0.365 0.307 0.347 0.311 0.337 0.329 0.343 0.328 0.346 0.316 0.365 0.349 0.380 0.310 0.350 0.295 0.338
Weather 0.255 0.284 0.262 0.300 0.277 0.305 0.275 0.286 0.287 0.281 0.264 0.273 - - 0.294 0.326 0.300 0.318 0.292 0.315 0.279 0.306
Average 0.335 0.365 0.348 0.378 0.361 0.386 0.361 0.371 0.357 0.361 0.372 0.373 0.407 0.428 0.402 0.440 0.452 0.420 0.457 0.409 0.414 0.396
1st Count 6 1 0 1 4 1 0 1 0 1 2
Short-Term Forecasting
To evaluate the effectiveness of the learned TSFM via our FeDaL in short-term forecasting tasks, we conduct experiments on the M4 dataset, following the protocols of zhou2023one . As shown in Table 4, FeDaL significantly outperforms baselines. While it performs slightly below FFTS on SMAPE by a narrow margin of 0.07%, FeDaL reduces MASE by 2% to 38% and achieves 5% to 22% improvements in SMAPE and OWA. These further confirm the ability of FeDaL to learn cross-domain representations that generalize effectively in short-term forecasting tasks.
Table 4: Average short-term forecasting results on M4 dataset. Bold: best; Underline: second best. None of these datasets were included in pretraining. * denotes a “former“ suffix. Full results are provide in Table˜16.
Models
FeDaL
FFTS
FedAvg
MOMENT
Time-LLM
GPT4TS
TimesNet
PatchTST
N-HiTS
N-BEATS
ETS.*
LightTS
DLinear
FED.
Stationary
Auto.*
Average
SMAPE 11.412 11.404 12.342 14.593 11.983 12.69 12.88 12.059 12.035 12.25 14.718 13.525 13.639 13.16 12.78 12.909
MASE 1.489 1.522 1.753 2.161 1.595 1.808 1.836 1.623 1.625 1.698 2.408 2.111 2.095 1.775 1.756 1.771
OWA 0.818 0.831 0.926 1.103 0.859 0.94 0.955 0.869 0.869 0.896 1.172 1.051 1.051 0.949 0.93 0.939
Imputation
Imputation aims to reconstruct missing values in time series based on partially observed data. We evaluate our method on five widely used real-world datasets, including four ETT datasets, and Weather, where the data-missing is common. Following the evaluation protocol of GPT4TS zhou2023one , we randomly mask time points at varying ratios to simulate different levels of corruption. As shown in Table˜5, the TSFM trained via our FeDaL consistently outperforms all baselines across datasets. Compared to the state-of-the-art centralized model GPT4TS, FeDaL reduces the average MSE by 22.84%. When compared to FL baselines FFTS and FedAvg, FeDaL achieves 12.64% and 27.62% relative MSE reductions, respectively. These demonstrate the effectiveness of FeDaL in learning common temporal patterns that generalize well even under missing-data conditions.
Table 5: Average imputation performance for randomly masked time series (input length 96), evaluated across four mask ratios $\{12.5\%,25\%,37.5\%,50\%\}$. Bold: best; Underline: second best. Full results in Table˜18.
Models FeDaL FFTS FedAvg TimeMixer GPT4TS PatchTST TimesNet DLinear Fed. Auto. Stationary LightTS In.
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.022 0.090 0.024 0.090 0.034 0.120 0.036 0.123 0.028 0.105 0.047 0.14 0.12 0.253 0.027 0.107 0.104 0.218 0.093 0.206 0.062 0.177 0.051 0.15 0.071 0.188
ETTh2 0.018 0.071 0.017 0.074 0.026 0.098 0.028 0.102 0.021 0.084 0.029 0.102 0.208 0.327 0.022 0.088 0.046 0.151 0.096 0.208 0.101 0.215 0.029 0.105 0.156 0.292
ETTm1 0.054 0.147 0.058 0.160 0.054 0.154 0.073 0.192 0.069 0.173 0.115 0.224 0.202 0.329 0.078 0.187 0.284 0.373 0.201 0.306 0.117 0.246 0.103 0.214 0.161 0.279
ETTm2 0.034 0.106 0.046 0.135 0.062 0.154 0.038 0.120 0.048 0.141 0.065 0.163 0.367 0.436 0.049 0.146 0.119 0.25 0.142 0.259 0.163 0.279 0.055 0.156 0.337 0.452
Weather 0.024 0.048 0.029 0.059 0.034 0.050 0.039 0.083 0.031 0.056 0.03 0.054 0.076 0.171 0.03 0.054 0.055 0.117 0.052 0.11 0.099 0.203 0.031 0.057 0.045 0.104
Average 0.030 0.092 0.035 0.104 0.042 0.115 0.043 0.124 0.039 0.112 0.057 0.137 0.195 0.303 0.041 0.116 0.122 0.222 0.117 0.218 0.108 0.224 0.054 0.136 0.154 0.263
1st Count 11 2 0 0 0 0 0 0 0 0 0 0 0
Anomaly Detection
Anomaly detection focuses on identifying abnormal patterns or segments. We evaluate FeDaL on five widely used datasets, including SMD, MSL, SMAP, SwaT, and PSM, following the evaluation protocols as GPT4TS zhou2023one to ensure fair comparison. The results as shown in Table˜6, the FeDaL-trained TSFM achieves the best overall performance across all datasets. It consistently outperforms prior state-of-the-art models, including ModernTCN and the general-purpose TSFM Moment, by 2.40% and 5.17%, respectively. Notably, from an FL perspective, FeDaL also delivers significant gains over FL baselines such as FedAvg and FFTS by 0.96% and 3.69%, respectively. These demonstrate FeDaL’s strong capability to capture global temporal invariances and its superior generalization to complex anomaly detection across heterogeneous client distributions.
Table 6: Anomaly detection results. We calculate the F1-score (%) for each dataset and statics the average F1-score. Bold: the best, Underline: the second best. * denotes a “former“ suffix. Full results are in Table˜19.
Model
FeDaL
FFTS
FedAvg
Moment
GPT4TS
MTCN
TimesNet
FED.*
LightTS
DLinear
NS.*
Auto.*
Pyra.*
Anomaly
In.*
Re.*
SMD
88.46
89.88
88.44
84.94
86.89
85.81
85.81
85.08
82.53
77.10
84.62
85.11
83.04
85.49
81.64
75.32
MSL
89.05
88.42
82.32
81.45
82.45
84.92
85.15
78.57
78.95
84.88
77.50
79.05
84.86
83.31
84.06
84.40
SMAP
71.70
71.38
70.78
69.43
72.88
71.26
71.52
70.76
69.21
69.26
71.09
71.12
71.09
71.18
69.92
70.40
SwaT
95.40
91.12
90.23
91.90
94.23
93.86
91.74
93.19
93.33
87.52
79.88
92.74
91.78
83.10
81.43
82.80
PSM
98.88
98.54
95.86
93.96
97.13
97.23
97.47
97.23
97.15
93.55
97.29
93.29
82.08
79.40
77.10
73.61
Avg. F1
88.70
87.86
85.54
84.34
86.72
86.62
86.34
84.97
84.23
82.46
82.08
84.26
82.57
80.50
78.83
77.31
Classification
To assess cross-task generalization, we evaluate the learned TSFM on time series classification using 10 UEA bagnall2018uea and 91 UCR dau2019ucr subsets spanning diverse domains. We adopt Linear Probing method by attaching a linear classifier to the frozen TSFM, directly measuring the quality of learned representations. As shown in Figure˜5, our FeDaL-trained TSFM consistently outperforms all baselines, including task-specific models, GPT4TS (fine-tuned), and Moment (generalized). Notably, FeDaL surpasses FL baselines such as FedAvg and FFTS by 6.4%/6.1% on UEA and 9.5%/3.2% on UCR, highlighting its ability to learn domain-invariant features and mitigate cross-client biases.
Figure 5: Classification results. Best viewed in color and with zoom.
4.3 Federated Scaling Behaviors
While prior work has examined scaling laws of centralized TSFMs, we investigate how federated pretraining scales with respect to (i) data size, (ii) number of clients, and (iii) client participation rate. Specifically, we vary: (i) data size from 40B to $\sim$231B with fixed 174 clients; (ii) client count from 30 to 174 under fixed total data; and (iii) participation rate from 10% to 100%. As shown in Figure˜6, larger data consistently improves performance, more clients yield better representations (even under fixed total data), and higher participation enhances aggregation and mitigates drift. These results indicate that federated TSFM pretraining benefits from scaling in data and client diversity, emphasizing coverage and participation over model size for improved generalization.
Figure 6: Scaling behaviors across tasks. Color codes: Blue – Avg. full/few-shot forecasting, Orange – Zero-shot forecasting, Green – Imputation, Red – Classification, Purple – Anomaly Detection. Y-axis for forecasting/imputation is inverted (lower is better). Full details and plots in Section˜C.5.
5 Conclusion
We propose Federated Dataset Learning (FeDaL), a brand-new framework for pretraining TSFMs under federated setting. By introducing Domain Bias Elimination (DBE) and Global Bias Elimination (GBE), Our FeDaL effectively mitigates domain bias caused by heterogeneous time series datasets, enabling the learning of domain-invariant temporal representations. Experiments on eight downstream tasks confirm FeDaL’s strong cross-domain generalization and superior performance. We also provide the first federated scaling analysis, offering empirical guidance for future TSFM design.
Appendix
This Appendix provides supplementary information and implementation details omitted from the main text, including:
  • More Related Work (Appendix˜A): A comprehensive review of relevant literature, covering time series foundation models, FL for heterogeneous data, and foundation models.
  • Implementation Details (Appendix˜B): Detailed descriptions of the techniques employed, training configurations, experimental setup—including benchmark procedures, data processing pipelines, and baseline methods.
  • Full Results (Appendix˜C): Complete presentation of all results discussed in the main text.
Appendix A More Related Work
Time Series Foundation Models
Pre-trained models with scalability can evolve into foundation models, characterized by increasing model capacity and pre-training scale, enabling them to address diverse data. Large Language Models (LLMs) have demonstrated advanced capabilities, including in-context learning and few/zero-shot learning [34]. However, due to the heterogeneity of time series data, the development of GPT-style time series foundation models (TSFMs) has been limited. Research on TSFMs is in its early stages, with existing efforts categorized into two groups, one of which is LLM-empowered time series analysis. For instance, FPT [82] utilizes GPT-2 as a representation extractor for time series, fine-tuning it on downstream datasets and tasks. LLM4TS [7] encodes time series into numerical tokens for LLMs, demonstrating model scalability in forecasting tasks. Time-LLM [34] explores prompting techniques to enhance prediction, inspiring cross-modality representation ability through language templates. UniTime [46] introduces cross-domain techniques to fine-tune pre-trained LLMs, reducing prediction biases. Nevertheless, the performance of these LLM-based approaches is heavily dependent on the LLM backbone and cross-modal design. In contrast, our proposed model is pre-trained natively on time series, eliminating the need for extra modality alignment. Another category of models involves pre-training on large-scale time series datasets. For instance, ForecastFPN [23] is trained on synthetic series to achieve zero-shot forecasting. CloudOps [63] employs masked modeling on Transformers for domain-specific forecasting. TimeGPT-1 [27] introduces the first commercial API for zero-shot forecasting, marking a milestone in the field. PreDcT [18], a decoder-only Transformer pre-trained on Google Trends, demonstrates notable zero-shot capabilities. Furthermore, MOMENT [28], Moirai [64], TimeFM [19], Chronos [2], and Timer [48] pre-train on ultra-large time series datasets and release corresponding datasets, showcasing excellent cross-task performance empirically. This suggests that training TSFMs is a resource-intensive endeavor.
Federated Learning for Heterogeneous Data
Federated Learning (FL) [53] enables decentralized model training across distributed clients without sharing raw data. A key challenge in FL is statistical heterogeneity (non-IID data) [13], which often leads to performance degradation due to distributional shifts across clients. Early solutions reduce client-server divergence through alignment or regularization, but typically assume mild heterogeneity and shared representation spaces—assumptions that break down in highly diverse domains. To address this, Personalized Federated Learning (PFL) aims to tailor models to individual clients. Existing methods include regularization-based decomposition [30, 40], partial model sharing [42, 16], adaptive aggregation [76], and meta-learning strategies [25]. While effective in some settings, PFL generally targets image or text data and assumes transferable low-level features [9], which do not generalize well to time series due to deeper cross-domain disparities in resolution, semantics, and physical context. Moreover, PFL prioritizes personalized models over unified generalization, making it ill-suited for time series foundation model (TSFM) pretraining, where cross-domain generalization is essential. To fill this gap, we propose Federated Dataset Learning (FeDaL)—a unified framework designed for TSFM training in decentralized, heterogeneous environments. By introducing Domain Bias Elimination (DBE) and Global Bias Elimination (GBE), FeDaL explicitly mitigates domain-specific biases at both client and server levels, enabling robust, scalable learning of generalized temporal representations.
Federated Learning for Foundation Models
FL is emerging as a powerful paradigm for training and adapting foundation models [70, 8, 10], especially in scenarios where centralizing large-scale, heterogeneous data is impractical due to computational or privacy constraints [9, 83, 32]. Unlike centralized strategies that require massive data aggregation and uniform model scaling, FL distributes the workload across clients, allowing each to contribute without sharing raw data. This decentralization not only reduces the computational burden but also preserves privacy while naturally accommodating data heterogeneity [83, 32]. Recent advances have explored FL for fine-tuning LLMs in personalized settings [49, 71, 32], demonstrating its potential to adapt pretrained models to diverse user data. In the time series domain, FL has also been used to fine-tune or train models for domain-specific tasks [44]. However, training foundation models for time series from scratch in federated settings presents new opportunities: FL offers a natural framework to leverage cross-domain data without mixing them, avoiding undesirable artifacts from centralized blending and enabling more principled treatment of domain-specific structures [9]. Despite these advantages, cross-domain bias—a key obstacle to generalization—has received little attention in federated TSFM training. Our proposed FeDaL specifically designed to address this gap. By explicitly mitigating client- and domain-level biases, FeDaL enables the training of unified TSFMs that capture high-level temporal representations transferable across domains and tasks.
Appendix B Implementation Details
Decoder-only Transformer as Local Models
We employ the Transformer as the model backbone due to its excellent scalability. Inspired by the significant advancements in Decoder-only LLMs capable of iterative generation [27, 82], and recognizing the need for processing variable-length time series [48, 74], we adopt an auto-regressive approach for generative pre-training using standard Decoder-only Transformer architectures. The next-token prediction can be formulated as:
$P(U)=\prod_{i=1}^{N}p(u_{i}\mid u_{<i}),\quad\text{where }U=\{u_{1},u_{2},...,u_{N}\}.$ (9)
For the tokenization of a given input time series $\bm{X}$}, we employed a segment-wise tokenization strategy, representing $\bm{X}$} as $\{x_{1},x_{2},\ldots,x_{NS}\}$ with a unified context length $NS$. In this approach, a time series token is defined as a consecutive segment of length $S$, covering the series variations: $\mathbf{s}_{i}=\{x_{(i-1)S+1},\ldots,x_{iS}\}\in\mathbb{R}^{S}$, where $i=1,\ldots,N$. Subsequently, a time series segment is incorporated into the learnable position encoding, followed by the standard autoregressive Transformer update step. Utilizing the causal attention mechanism of the Decoder-only Transformer, the model autoregressively generates the subsequent segment $\mathbf{s}_{i+1}$ based on the previous segment $\mathbf{s}_{i}$. Consequently, generative pre-trained models are endowed with the flexibility to handle variable context lengths during inference and excel at multi-step generation through iterative sliding and enlarging of input tokens.
Fourier-based Perturbation for Privacy Preservation
To ensure privacy during core-set transmission, we introduce a Fourier-based perturbation mechanism that obfuscates sensitive raw temporal patterns while preserving high-level semantic structure. This builds on the observation that in time series, the phase of the Fourier transform captures semantic trends (e.g., shape, rhythm), while the amplitude encodes finer-grained details (e.g., scale, local fluctuations). Formally, given a core-set $\mathcal{C}$, we apply a discrete Fourier transform (DFT) to obtain its frequency-domain representation $\mathcal{F}(\mathcal{C})=A+iP$, where $A$ and $P$ denote amplitude and phase, respectively. We then perturb only the amplitude component:
$\mathcal{F}(\mathcal{C}_{\text{init}})=A+iP,$ (10)
where $A$ and $P$ denote the amplitude and phase components, respectively. We then perturb only the amplitude component with Gaussian noise:
$\tilde{\mathcal{C}}=\mathcal{F}^{-1}\left((A+\epsilon\cdot\mathcal{N}(0,1))+iP\right),$ (11)
where $\mathcal{F}^{-1}$ denotes the inverse Fourier transform, and $\epsilon$ controls the perturbation strength. This ensures that while the fine-grained raw sequence details are obfuscated, the high-level temporal semantics embedded in the phase remain intact. To ensure that $\mathcal{C}^{\prime}$ remains semantically faithful to the original data, we align its latent representation with that of the sampled mini-batch $\tilde{\mathcal{X}}$:
$\mathcal{L}_{\text{align}}({\mathcal{C}}^{\prime})=\frac{1}{K}\|\sum\nolimits_{c_{j}\in\mathcal{C}^{\prime}}f_{\theta}(c_{j})-\sum\nolimits_{x_{k}\in\tilde{\mathcal{X}}}f_{\theta}(x_{k})\|_{2}^{2}.$ (12)
The semantically aligned core-set $\mathcal{C}$, optimized from the perturbed $\mathcal{C}^{\prime}$ via Eq. 12, is then uploaded to the server for the server-side core-set tuning via the privacy-preserving manner.
T-SNE Visualization on Core-set
To further assess the semantic fidelity and privacy-preserving effect of our core-set pipeline, we visualize the t-SNE embeddings of three data subsets for several clients: (i) the original local mini-batch $\tilde{\mathcal{X}}$ used for core-set construction, (ii) the gradient-matched core-set $\mathcal{C}_{\text{init}}$, and (iii) the final perturbed and aligned core-set $\mathcal{C}^{\prime}$ sent to the server. As shown in Figure˜7, the initial core-set (green) exhibits tight overlap with the original batch (pink), confirming that $\mathcal{C}_{\text{init}}$ successfully approximates local gradients—but also indicating potential privacy risks if shared directly. In contrast, the perturbed and aligned core-set (blue) is visibly displaced, thanks to our Fourier-based amplitude perturbation. Nonetheless, $\mathcal{C}^{\prime}$ maintains global structure and cluster consistency with $\tilde{\mathcal{X}}$, highlighting that our representation alignment objective effectively preserves semantic integrity while masking raw data characteristics. These results validate our claim that the core-set tuning pipeline provides a strong utility-privacy trade-off in the server-side core-set tuning.
Figure 7: t-SNE visualization of core-set across clients. Each subfigure corresponds to a different client. The initial core-set closely aligns with the local data, while the perturbed core-set shows semantic similarity with added privacy-preserving shifts. Best viewed in color and zoom-in.
Time Series Decomposition
Time series decomposition aims to separate a sequence into interpretable components, typically trend, seasonal, and residual parts. This allows models to better capture structured temporal patterns and isolate noise or biases. In our framework, we follow the decomposition strategy proposed in [68], applying it directly in the latent representation space rather than the raw input domain. Given a latent representation sequence $\mathbf{h}=f_{\theta^{b}}({\bm{X}}_{i}^{\text{masked}})\in\mathbb{R}^{L\times d}$, we decompose it into trend and seasonal components using a moving average operator with period $\tau$:
$\mathbf{h}^{\text{trend}}=\texttt{MA}_{\tau}(\mathbf{h}),\quad\mathbf{h}^{\text{season}}=\mathbf{h}-\mathbf{h}^{\text{trend}}.$ (13)
Here, $\texttt{MA}_{\tau}(\cdot)$ denotes a $\tau$-point moving average operator. The resulting $\mathbf{H}^{\text{trend}}$ captures low-frequency temporal dynamics, while $\mathbf{H}^{\text{season}}$ reflects higher-frequency but still structured periodic variations. We omit the residual component to reduce complexity and avoid modeling unstructured noise. To estimate persistent client-specific deviations, we apply temporal averaging over both components:
$\mathbf{b}_{i}^{\text{trend}}=\texttt{Mean}(\mathbf{h}_{i}^{\text{trend}}),\quad\mathbf{b}_{i}^{\text{season}}=\texttt{Mean}(\mathbf{h}_{i}^{\text{season}}),$ (14)
which serve as compact and effective approximations of domain-specific biases accumulated from heterogeneous time series distributions.
Basic Training Configuration
For both federated representation learning and downstream adaptation, we set the batch size to 2048. The local training uses 10 epochs per round, with total global rounds of 200 for federated representation learning and 1000 for downstream adaptation. The initial learning rate is fixed at $1e^{-2}$. Additionally, we configure the core-set size as $K=1024$ for federated representation learning and $K=3072$ for downstream adaptation.
Benchmark Details
We evaluate FeDaL across two major settings: Federated Representation Learning and Downstream Adaptation. For Federated Representation Learning, we use two large-scale cross-domain time series datasets: (1) UTSD [^1] https://huggingface.co/datasets/thuml/UTSD, containing over 1 billion time points across 7 domains (e.g., energy, environment, health, IoT, nature); and (2) CTSD [^2] https://forecastingdata.org/, sampled from the Monash Time Series Forecasting Repository, comprising 500 million time points from 6 domains including ergonomics, transportation, health, energy, nature, and web. For Downstream Adaptation, we pretrain on the large-scale LOTSA dataset [64]  [^3] https://huggingface.co/datasets/Salesforce/lotsa_data, which contains 231 billion observations spanning multiple domains. We evaluate the pretrained TSFM from FeDaL on four downstream tasks: long- and short-term forecasting, imputation, classification, and anomaly detection. For long-term forecasting and imputation, we follow the GPT4TS [82] protocol, using ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Illness datasets (details in Table˜8). For short-term forecasting, we adopt the M4 dataset, again following GPT4TS, covering various temporal granularities (details in Table˜9). For classification, we follow the TimesBERT [74] setup, using 10 representative datasets from UEA and 91 from UCR to cover diverse domains (see Table˜10). For anomaly detection, we adopt the FFTS [9] benchmark, evaluating on SMD, MSL, SMAP, SWaT, and PSM datasets (details in Table˜11). Any datasets overlapping with LOTSA have been excluded from the pretraining phase.
Table 7: Dataset statistics about the CTSD dataset. The channels indicates the number of time series. ‘Min‘ and ‘Max‘ denote the shortest and longest sequence lengths, respectively, while ‘Fixed‘ signifies datasets with uniform sequence lengths.
Domain Dataset # Channels Frequency Length
Economic Bitcoin 18 Daily Min: 2659 Max: 4581
FRED-MD 107 Monthly Fixed: 728
NN5 111 Daily Fixed: 791
Transport Pedestrain Counts 66 Hourly Min: 576 Max: 96424
Rideshare 2304 Daily Fixed: 541
San Francisco Traffic 862 Hourly/Weekly Fixed: 17544
Health COVID Deaths 266 Daily Min: 212 Max: 212
Hospital 767 Monthly Fixed: 84
Energy London Smart Meters 5560 Daily Min: 288 Max: 39648
Wind Farms 339 Minutely Min: 6345 Max: 527040
Wind Power 1 Second Fixed: 7397147
Electricity 321 Hourly/Weekly Fixed: 26304
Nature KDD Cup 2018 270 Hourly Min: 9504 Max: 10920
Oikolab Weather 3010 Daily Min: 1332 Max: 65981
Temperature Rain 32072 Daily Fixed: 725
Web Web Traffic 145063 Daily Fixed: 803
Table 8: Dataset statistics about long-term forecasting (imputation) dataset. The channels indicates the number of time series (i.e., variables), and the size is organized in (training, validation, testing). Note that Illness dataset is not included in the imputation task.
Dataset Domain # Channels Frequency Size Forecast Length
ETTh1 Power 7 Hourly (8545, 2881, 2881) $\{96,192,336,720\}$
ETTh2 Power 7 Hourly (8545, 2881, 2881) $\{96,192,336,720\}$
ETTm1 Power 7 15 Minute (34465, 11521, 11521) $\{96,192,336,720\}$
ETTm2 Power 7 15 Minute (34465, 11521, 11521) $\{96,192,336,720\}$
Illness Epidemiology 7 Weekly (617, 74, 170) $\{24,36,48,60\}$
Weather Weather 21 10 Minute (36792, 5271, 10540) $\{96,192,336,720\}$
Table 9: Dataset statistics about short-term forecasting dataset. The channels indicates the number of time series (i.e., variables), and the size is organized in (training, validation, testing).
Dataset Domain # Channels Frequency Size Forecast Length
M4-Yearly Demographic 1 Yearly (23000, 0, 23000) 6
M4-Quarterly Finance 1 Quarterly (24000, 0, 24000) 8
M4-Monthly Industry 1 Monthly (48000, 0, 48000) 18
M4-Weekly Macro 1 Weekly (359, 0, 359) 13
M4-Daily Macro 1 Daily (4227, 0, 4227) 14
M4-Hourly Other 1 Hourly (414, 0, 414) 48
Table 10: Dataset statistics about classification dataset. The channels indicates the number of time series (i.e., variables), and the size is organized in (training, validation, testing).
Dataset # Channels Series Length Dataset Size Information (Frequency)
EthanolConcentration 3 1751 (261, 0, 263) Alcohol Industry
FaceDetection 144 62 (5890, 0, 3524) Face (250Hz)
Handwriting 3 152 (150, 0, 850) Handwriting
Heartbeat 61 405 (204, 0, 205) Heart Beat
JapaneseVowels 12 29 (270, 0, 370) Voice
PEMS-SF 963 144 (267, 0, 173) Transportation (Daily)
SelfRegulationSCP1 6 896 (268, 0, 293) Health (256Hz)
SelfRegulationSCP2 7 1152 (200, 0, 180) Health (256Hz)
SpokenArabicDigits 13 93 (6599, 0, 2199) Voice (11025Hz)
UWaveGestureLibrary 3 315 (120, 0, 320) Gesture
UCR Archive 1 * (*, 0, *) *
Table 11: Dataset statistics about anomaly detection dataset. The channels indicates the number of time series (i.e., variables), and the size is organized in (training, validation, testing).
Dataset # Channels Series Length Dataset Size Information (Frequency)
SMD 38 40 (566724, 141681, 708420) Server Machine
MSL 55 40 (44653, 11664, 73729) Spacecraft
SMAP 25 40 (108146, 27037, 427617) Spacecraft
SWaT 51 40 (396000, 99000, 449919) Infrastructure
PSM 25 40 (105984, 26497, 87841) Server Machine
Figure 8: Comparison between Domain-Mixed (DM) and Domain-Independent (DI) settings. In DM, each client contains time series from multiple domains. In contrast, in DI, each client contains time series from a single domain, though the same domain may appear across different clients with non-overlapping sequences.
Baseline
Table 12: Baseline information pertains to federated representation learning and its applications in forecasting, imputation, anomaly detection, and classification for downstream adaptation. Regarding the Foundation Model class, all are employed exclusively for the zero-shot long-term forecasting task, except for Moment, which is applied across the entire scenario.
Baseline Venue Task
Representation Forecasting Imputation Anomaly Classification
FL Methods
FedAvg AISTATS’17 [53]
FedProx MLSys’20 [41]
FedPer arXiv’19 [3]
FedRep ICML’21 [16]
FFTS AAAI’25 [9]
FeDaL (Ours) This paper
Deep Models and Traditional Machine Learning Methods
TimeMixer [61] ICLR’24 [61]
GPT4TS NeurIPS’23 [82]
PatchTST ICLR’23 [54]
TimesNet ICLR’23 [66]
DLinear AAAI’23 [73]
Fedformer ICML’22
Autoformer NeurIPS’21
Stationary NeurIPS’22 [47]
LightTS arXiv’22 [78]
Informer AAAI’21 [80]
Reformer ICLR’20 [36]
Time-LLM ICLR’24 [34]
N-HiTS AAAI’23 [6]
N-BEATS ICLR’20 [55]
ETSformer arXiv’22 [65]
MTCN ICLR’24 [51]
Pyraformer ICLR’22 [45]
Anomaly ICLR’22 [69]
DTW ICML’17 [17]
TS2Vec AAAI’22 [72]
T-Loss NeurIPS’19 [26]
TNC ICLR’21 [59]
TS-TCC IJCAI’21 [24]
TST /
CNN /
Encoder /
FCN /
MCDNN CVPR’12 [15]
ResNet CVPR’16 [31]
t-LeNet ECML’16 [38]
TWIESN TR’06 [50]
MLP /
XGBoost KDD’16 [14]
Rocket DMKD’20 [21]
LSTM NC’97 [33]
LSTNet SIGIR’18 [37]
LSSL ICLR’22 [29]
Transformer NeurIPS’17 [60]
Flowformer ICML’22 [67]
LogTransformer NeurIPS’19 [39]
TCN NeurIPS’19 [26]
Foundation Model
Moment ICML’24 [28]
Moiraismall ICML’24 [64]
Morialbase ICML’24 [64]
Moriallarge ICML’24 [64]
TimesFM ICML’24 [19]
Chronossmall TMLR’24 [2]
Chronosbase TMLR’24 [2]
Chronoslarge TMLR’24 [2]
We compare our proposed FeDaL against 54 different baselines in our experiments to demonstrate its effectiveness and superiority in federated time series representation learning and downstream time series analysis tasks. In this section, we introduce some of these representative baselines from both federated learning and deep time series modeling. Full information of baselines can be found in Table˜12. The key federated learning baselines include:
  • FedAvg [53]: A decentralized approach that enables devices to collaboratively learn a shared model by aggregating locally-computed updates, allowing for the training of deep networks on private and large datasets while reducing communication costs.
  • FedProx [41]: A federated learning framework that generalizes and re-parameterizes FedAvg to tackle systems and statistical heterogeneity, providing convergence guarantees and demonstrating more robust and accurate convergence behavior.
  • FedPer [3]: A federated learning baseline that personalizes models by keeping client-specific layers local while training shared base layers collaboratively.
  • FedRep [16]: A federated learning baseline that decouples learning into global feature representation (shared) and client-specific heads (local) for better personalization.
  • FFTS [9] A FL-based decentralized framework for training time series foundation models from scratch, which incorporates a Mixture-of-Experts mechanism on each client and employs a heterogeneous knowledge alignment strategy to address time series heterogeneity.
Details of deep time series model baseline is as follows:
  • LogTransformer [39]: A modified Transformer architecture that addresses the locality-agnostics and memory bottleneck issues in time series forecasting by incorporating convolutional self-attention and log-sparse. attention, achieving improved forecasting accuracy with reduced memory cost.
  • N-BEATS [55]: A deep neural architecture that achieves state-of-the-art performance in univariate time series point forecasting, using a stack of fully-connected layers with backward and forward residual links.
  • Reformer [36]: This model improves Transformer by using locality-sensitive hashing for attention and reversible residual layers. It offers better memory efficiency and speed for lengthy sequences without sacrificing performance.
  • Informer [80]: An optimized Transformer-based model for long-range time series prediction. It uses ProbSparse self-attention for efficiency, processes long inputs effectively, and employs a fast prediction decoder.
  • LightTS [78]: A lightweight MLP structure. It utilizes two downsampling strategies—spaced and sequential sampling—on the MLP structure, capitalizing on the fact that downsampled time series generally maintain most of their original information.
  • ETSformer [65]: A Transformer architecture that leverages the principle of exponential smoothing to improve traditional Transformers for time-series forecasting, offering better decomposition capability, interpretability, and long-term forecasting efficiency.
  • Stationary (NS*) [47]: A framework that addresses the over-stationarization problem in time series forecasting by combining Series Stationarization and De-stationary Attention modules, which unify statistics for better predictability while recovering intrinsic non-stationary information.
  • Autoformer [68]: A decomposition architecture that leverages leveraging an Auto-Correlation mechanism for long-term forecasting, which efficiently discovers dependencies and aggregates representations at the sub-series level.
  • FEDformer [81]: A forecasting method that combines seasonal-trend decomposition with a frequency-enhanced Transformer, capturing both the global profile and detailed structures of time series.
  • Pyraformer [45]: It features hierarchical pyramidal attention modules with binary trees to capture temporal dependencies across different ranges efficiently, both in time and memory complexity.
  • AnomalyTransformer [69]: A approach for unsupervised time series anomaly detection that leverages the self-attention mechanism to compute association discrepancy, which captures the adjacent-concentration bias of anomalies.
  • TimesNet [66]: A framework that transforms 1D time series into 2D tensors to model complex temporal variations, leveraging the multi-periodicity of time series to adaptively discover and extract intraperiod- and interperiod-variations.
  • PatchTST [54]: This method divides the time series into patches at the sub-series level for input to the Transformer. Each channel holds a univariate time series, sharing the same embedding and Transformer weights across all series.
  • DLinear [73]: DLinear integrates decomposition schemes from Autoformer and FEDformer with linear layers to model time series data tables. It effectively summarizes trend and seasonal components, enhancing performance on datasets rich in trends.
  • N-HiTS [6]: A forecasting model that addresses the challenges of long-horizon forecasting by incorporating hierarchical interpolation and multi-rate data sampling techniques.
  • GPT4TS [82]: This model is designed for time series analysis across various scenarios, achieved by fine-tuning a pre-trained language model, specifically GPT2, for the time series domain.
  • Time-LLM [34]: A reprogramming framework that repurposes large language models (LLMs) for general time series forecasting by aligning time series data with natural language modalities through text prototypes and Prompt-as-Prefix (PaP) techniques.
  • TimeMixer [61]: An advanced time series forecasting baseline that employs a multi-scale mixing architecture with Past Decomposition Mixing (PDM) and Future Multi-predictor Mixing (FMM) modules to effectively disentangle and integrate seasonal and trend patterns.
Evaluation Metrics
For evaluation metrics in forecasting and imputation tasks, we utilize the mean square error (MSE) and mean absolute error (MAE) for long-term forecasting. In terms of the short-term forecasting on M4 benchmark, we adopt the symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA) as in N-BEATS. Note that OWA is a specific metric utilized in the M4 competition. The calculations of these metrics are as follows:
MSE $=\frac{1}{H}\sum_{h=1}^{T}(\mathbf{Y}_{h}-\hat{\mathbf{Y}}_{h})^{2},$ (15)
MAE $=\frac{1}{H}\sum_{h=1}^{H}|\mathbf{Y}_{h}-\hat{\mathbf{Y}}_{h}|,$
SMAPE $=\frac{200}{H}\sum_{h=1}^{H}\frac{|\mathbf{Y}_{h}-\hat{\mathbf{Y}}_{h}|}{|\mathbf{Y}_{h}|+|\hat{\mathbf{Y}}_{h}|},$
MAPE $=\frac{100}{H}\sum_{h=1}^{H}\frac{|\mathbf{Y}_{h}-\hat{\mathbf{Y}}_{h}|}{|\mathbf{Y}_{h}|},$
MASE $=\frac{1}{H}\sum_{h=1}^{H}\frac{|\mathbf{Y}_{h}-\hat{\mathbf{Y}}_{h}|}{\frac{1}{H-s}\sum_{j=s+1}^{H}|\mathbf{Y}_{j}-\mathbf{Y}_{j-s}|},$
OWA $=\frac{1}{2}\left[\frac{\text{SMAPE}}{\text{SMAPE}_{\textrm{Na\"{\i}ve2}}}+\frac{\text{MASE}}{\text{MASE}_{\textrm{Na\"{\i}ve2}}}\right],$
where $s$ is the periodicity of the time series data. $H$ denotes the number of data points (i.e., prediction horizon in our cases). $\mathbf{Y}_{h}$ and $\hat{\mathbf{Y}}_{h}$ are the $h$-th ground truth and prediction where $h\in\{1,\cdots,H\}$.
In addition, we used Precision (P), Recall (R), and F1-Score (F1) to simply quantify the performance of our FeDaL and baselines on the anomaly detection task, these can be formulated as:
$\text{P}=\frac{\text{TP}}{\text{TP}+\text{FP}},$ (16)
$\text{R}=\frac{\text{TP}}{\text{TP}+\text{FN}},$
$\text{F1}=2\times\frac{\text{P}\times\text{R}}{\text{P}+\text{R}},$
where TP (True Positives), FP (False Positives), and FN (False Negatives) represent the number of samples correctly labeled as anomalous, the number of samples incorrectly labeled as anomalous, and the number of samples that were not labeled as anomalous by the model but were actually anomalous, respectively. For classification tasks, we used Accuracy to evaluate the performance.
Appendix C Full Results
This section presents the complete experimental results across full-shot and zero-shot long-term forecasting, short-term forecasting, imputation, anomaly detection, and classification. It also includes detailed plots and analysis of federated scaling behaviors during time series foundation model training.
C.1 Long-term Forecasting
The full long-term forecasting results are provided in Table˜13 and Table˜14. Table˜13 demonstrates that FeDaL-trained TSFM consistently outperforms state-of-the-art deep time series models specifically designed for long-term forecasting. Furthermore, Table˜14 shows that under zero-shot evaluation, our approach surpasses both advanced centralized pretrained TSFMs and federated pretraining baselines, highlighting its strong generalization across domains and tasks.
Table 13: Full long-term forecasting results comparing our proposed FeDaL with advanced deep time series models. Bold: the best, Underline: the second best.
Dataset Method FeDaL (Ours) TimeMixer Time-LLM GPT4TS DLinear PatchTST TimesNet FED. Auto. Stationary ETS. LightTS Informer Re.
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.340 0.371 0.373 0.401 0.362 0.392 0.376 0.397 0.375 0.399 0.370 0.399 0.384 0.402 0.376 0.419 0.449 0.459 0.513 0.491 0.494 0.479 0.424 0.432 0.865 0.713 0.837 0.728
192 0.371 0.400 0.436 0.430 0.398 0.418 0.416 0.418 0.405 0.416 0.413 0.421 0.436 0.429 0.420 0.448 0.500 0.482 0.534 0.504 0.538 0.504 0.475 0.462 1.008 0.792 0.923 0.766
336 0.392 0.420 0.484 0.458 0.430 0.427 0.442 0.433 0.439 0.443 0.422 0.436 0.491 0.469 0.459 0.465 0.521 0.496 0.588 0.535 0.574 0.521 0.518 0.488 1.107 0.809 1.097 0.835
720 0.417 0.445 0.497 0.482 0.442 0.457 0.477 0.456 0.472 0.490 0.447 0.466 0.521 0.500 0.506 0.507 0.514 0.512 0.643 0.616 0.562 0.535 0.547 0.533 1.181 0.865 1.257 0.889
Avg. 0.380 0.409 0.448 0.443 0.408 0.423 0.465 0.455 0.422 0.437 0.413 0.430 0.458 0.450 0.440 0.460 0.496 0.487 0.570 0.537 0.542 0.510 0.491 0.479 1.040 0.795 1.029 0.805
ETTh2 96 0.267 0.325 0.289 0.340 0.268 0.328 0.285 0.342 0.289 0.353 0.274 0.336 0.340 0.374 0.358 0.397 0.346 0.388 0.476 0.458 0.340 0.391 0.397 0.437 3.755 1.525 2.626 1.317
192 0.333 0.369 0.370 0.389 0.329 0.375 0.354 0.389 0.383 0.418 0.339 0.379 0.402 0.414 0.429 0.439 0.456 0.452 0.512 0.493 0.430 0.439 0.520 0.504 5.602 1.931 11.12 2.979
336 0.359 0.392 0.386 0.413 0.368 0.409 0.373 0.407 0.448 0.465 0.329 0.380 0.452 0.452 0.496 0.487 0.482 0.486 0.552 0.551 0.485 0.479 0.626 0.559 4.721 1.835 9.323 2.769
720 0.377 0.420 0.412 0.432 0.372 0.420 0.406 0.441 0.605 0.551 0.379 0.422 0.462 0.468 0.463 0.474 0.515 0.511 0.562 0.560 0.500 0.497 0.863 0.672 3.647 1.625 3.874 1.697
Avg 0.334 0.377 0.364 0.394 0.334 0.383 0.381 0.412 0.431 0.446 0.330 0.379 0.414 0.427 0.437 0.449 0.450 0.459 0.526 0.516 0.439 0.452 0.602 0.543 4.431 1.729 6.736 2.191
ETTm1 96 0.258 0.321 0.320 0.357 0.272 0.334 0.292 0.346 0.299 0.343 0.290 0.342 0.338 0.375 0.379 0.419 0.505 0.475 0.386 0.398 0.375 0.398 0.374 0.400 0.672 0.571 0.538 0.528
192 0.297 0.343 0.361 0.380 0.310 0.358 0.332 0.372 0.335 0.365 0.332 0.369 0.374 0.387 0.426 0.441 0.553 0.496 0.459 0.444 0.408 0.410 0.400 0.407 0.795 0.669 0.658 0.592
336 0.327 0.381 0.392 0.404 0.352 0.384 0.366 0.394 0.369 0.386 0.366 0.392 0.410 0.411 0.445 0.459 0.621 0.537 0.495 0.464 0.435 0.428 0.438 0.438 1.212 0.871 0.898 0.721
720 0.392 0.415 0.452 0.440 0.383 0.411 0.417 0.421 0.425 0.421 0.416 0.420 0.478 0.450 0.543 0.490 0.671 0.561 0.585 0.516 0.499 0.462 0.527 0.502 1.166 0.823 1.102 0.841
Avg 0.319 0.365 0.381 0.395 0.329 0.372 0.388 0.403 0.357 0.378 0.351 0.380 0.400 0.406 0.448 0.452 0.588 0.517 0.481 0.456 0.429 0.425 0.435 0.437 0.961 0.734 0.799 0.671
ETTm2 96 0.157 0.240 0.175 0.258 0.161 0.253 0.173 0.262 0.167 0.269 0.165 0.255 0.187 0.267 0.203 0.287 0.255 0.339 0.192 0.274 0.189 0.280 0.209 0.308 0.365 0.453 0.658 0.619
192 0.220 0.292 0.235 0.298 0.219 0.293 0.229 0.301 0.224 0.303 0.220 0.292 0.249 0.309 0.269 0.328 0.281 0.340 0.280 0.339 0.253 0.319 0.311 0.382 0.533 0.563 1.078 0.827
336 0.267 0.342 0.298 0.341 0.271 0.329 0.286 0.341 0.281 0.342 0.274 0.329 0.321 0.351 0.325 0.366 0.339 0.372 0.334 0.361 0.314 0.357 0.442 0.466 1.363 0.887 1.549 0.972
720 0.399 0.400 0.391 0.395 0.352 0.379 0.378 0.401 0.397 0.421 0.362 0.385 0.408 0.403 0.421 0.415 0.433 0.432 0.417 0.413 0.414 0.413 0.675 0.587 3.379 1.338 2.631 1.242
Avg 0.261 0.319 0.275 0.323 0.251 0.313 0.284 0.339 0.267 0.333 0.255 0.315 0.291 0.333 0.305 0.349 0.327 0.371 0.306 0.347 0.293 0.342 0.409 0.436 1.410 0.810 1.479 0.915
Weather 96 0.139 0.199 0.162 0.209 0.147 0.201 0.162 0.212 0.176 0.237 0.149 0.198 0.172 0.220 0.217 0.296 0.266 0.336 0.173 0.223 0.197 0.281 0.182 0.242 0.300 0.384 0.689 0.596
192 0.174 0.230 0.208 0.250 0.189 0.234 0.204 0.248 0.220 0.282 0.194 0.241 0.219 0.261 0.276 0.336 0.307 0.367 0.245 0.285 0.237 0.312 0.227 0.287 0.598 0.544 0.752 0.638
336 0.240 0.279 0.252 0.287 0.262 0.279 0.254 0.286 0.265 0.319 0.245 0.282 0.280 0.306 0.339 0.380 0.359 0.395 0.321 0.338 0.298 0.353 0.282 0.334 0.578 0.523 0.639 0.596
720 0.300 0.310 0.340 0.343 0.304 0.316 0.326 0.337 0.333 0.362 0.314 0.334 0.365 0.359 0.403 0.428 0.419 0.428 0.414 0.410 0.352 0.288 0.352 0.386 1.059 0.741 1.130 0.792
Avg 0.213 0.255 0.241 0.272 0.225 0.257 0.237 0.270 0.248 0.300 0.225 0.264 0.259 0.287 0.309 0.360 0.338 0.382 0.288 0.314 0.271 0.334 0.261 0.312 0.634 0.548 0.803 0.656
ILI 24 1.265 0.714 1.979 0.860 1.285 0.727 2.063 0.881 2.215 1.081 1.319 0.754 2.317 0.934 3.228 1.260 3.483 1.287 2.294 0.945 2.527 1.020 8.313 2.144 5.764 1.677 4.400 1.382
36 1.329 0.800 1.893 0.862 1.404 0.814 1.868 0.892 1.963 0.963 1.430 0.834 1.972 0.920 2.679 1.080 3.103 1.148 1.825 0.848 2.615 1.007 6.631 1.902 4.755 1.467 4.783 1.448
48 1.409 0.768 2.129 0.936 1.523 0.807 1.790 0.884 2.130 1.024 1.553 0.815 2.238 0.940 2.622 1.078 2.669 1.085 2.010 0.900 2.359 0.972 7.299 1.982 4.763 1.469 4.832 1.465
60 1.418 0.810 2.155 0.938 1.531 0.854 1.979 0.957 2.368 1.096 1.470 0.788 2.027 0.928 2.857 1.157 2.770 1.125 2.178 0.963 2.487 1.016 7.283 1.985 5.264 1.564 4.882 1.483
Avg 1.355 0.773 2.039 0.899 1.435 0.801 1.925 0.903 2.169 1.041 1.443 0.797 2.139 0.931 2.847 1.144 3.006 1.161 2.077 0.914 2.497 1.004 7.382 2.003 5.137 1.544 4.724 1.445
$1^{\text{st}}$ Count
46 0 13 0 0 5 0 0 0 0 0 0 0 3
Table 14: Full results of zero-shot forecasting experiments. A lower MSE or MAE indicates a better prediction. TimesFM, due to its use of Weather datasets in pretraining, is not evaluated on this dataset and is denoted by a dash (-). Bold: the best, Underline: the second best.
Models
Federated Learning Methods Pretrained Time Series Foundation Models
FeDaL (Ours) FFTS FedAvg Moiraismall Moiraibase Moirailarge TimesFM Moment Chronossmall Chronosbase Chronoslarge
Metrics
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.347 0.381 0.344 0.382 0.412 0.409 0.401 0.402 0.376 0.392 0.349 0.379 0.414 0.404 0.688 0.557 0.466 0.409 0.440 0.393 0.441 0.390
192 0.398 0.410 0.395 0.410 0.420 0.421 0.388 0.412 0.412 0.413 0.434 0.415 0.465 0.434 0.688 0.560 0.530 0.450 0.492 0.426 0.502 0.424
336 0.425 0.452 0.438 0.445 0.440 0.444 0.433 0.428 0.433 0.428 0.495 0.445 0.503 0.456 0.675 0.563 0.570 0.486 0.550 0.462 0.576 0.467
720 0.457 0.469 0.445 0.457 0.480 0.522 0.439 0.454 0.447 0.444 0.611 0.510 0.511 0.481 0.683 0.585 0.615 0.543 0.882 0.591 0.835 0.583
Avg. 0.407 0.429 0.425 0.437 0.438 0.449 0.428 0.427 0.417 0.419 0.480 0.439 0.473 0.443 0.683 0.566 0.545 0.472 0.591 0.468 0.588 0.466
ETTh2 96 0.307 0.355 0.325 0.332 0.340 0.350 0.297 0.336 0.294 0.330 0.296 0.330 0.315 0.349 0.342 0.396 0.307 0.356 0.308 0.343 0.320 0.345
192 0.349 0.372 0.355 0.359 0.378 0.388 0.368 0.381 0.365 0.375 0.361 0.371 0.388 0.395 0.354 0.402 0.376 0.401 0.384 0.392 0.406 0.399
336 0.387 0.395 0.391 0.393 0.415 0.421 0.370 0.393 0.376 0.390 0.390 0.390 0.422 0.427 0.356 0.407 0.408 0.431 0.429 0.430 0.492 0.453
720 0.401 0.406 0.409 0.412 0.427 0.441 0.411 0.426 0.416 0.433 0.423 0.418 0.443 0.454 0.395 0.434 0.604 0.533 0.501 0.477 0.603 0.511
Avg. 0.361 0.382 0.370 0.374 0.390 0.401 0.361 0.384 0.362 0.382 0.367 0.377 0.392 0.406 0.361 0.409 0.424 0.430 0.405 0.410 0.455 0.427
ETTm1 96 0.289 0.346 0.307 0.352 0.303 0.371 0.418 0.392 0.363 0.356 0.380 0.361 0.361 0.370 0.654 0.527 0.511 0.423 0.454 0.408 0.457 0.403
192 0.317 0.369 0.324 0.392 0.344 0.384 0.431 0.405 0.388 0.375 0.412 0.383 0.414 0.405 0.662 0.532 0.618 0.485 0.567 0.477 0.530 0.450
336 0.370 0.420 0.399 0.438 0.426 0.432 0.433 0.412 0.416 0.392 0.436 0.400 0.445 0.429 0.672 0.537 0.683 0.524 0.662 0.525 0.577 0.481
720 0.464 0.426 0.426 0.498 0.438 0.453 0.462 0.432 0.460 0.418 0.462 0.420 0.512 0.471 0.692 0.551 0.748 0.566 0.900 0.591 0.660 0.526
Avg. 0.360 0.390 0.364 0.420 0.378 0.410 0.436 0.410 0.406 0.385 0.422 0.391 0.433 0.418 0.670 0.536 0.640 0.499 0.645 0.500 0.555 0.465
ETTm2 96 0.207 0.283 0.222 0.278 0.219 0.289 0.214 0.288 0.205 0.273 0.211 0.274 0.202 0.270 0.260 0.335 0.209 0.291 0.199 0.274 0.197 0.271
192 0.248 0.333 0.273 0.320 0.270 0.333 0.284 0.332 0.275 0.316 0.281 0.318 0.289 0.321 0.289 0.350 0.280 0.341 0.261 0.322 0.254 0.314
336 0.316 0.340 0.320 0.327 0.321 0.340 0.331 0.362 0.329 0.350 0.341 0.355 0.360 0.366 0.324 0.369 0.354 0.390 0.326 0.366 0.313 0.353
720 0.397 0.408 0.453 0.523 0.478 0.498 0.402 0.408 0.437 0.411 0.485 0.428 0.462 0.430 0.394 0.409 0.553 0.499 0.455 0.439 0.416 0.415
Avg. 0.292 0.341 0.317 0.362 0.322 0.365 0.307 0.347 0.311 0.337 0.329 0.343 0.328 0.346 0.316 0.365 0.349 0.380 0.310 0.350 0.295 0.338
Weather 96 0.159 0.212 0.172 0.218 0.201 0.207 0.198 0.222 0.220 0.217 0.199 0.211 - - 0.243 0.255 0.211 0.243 0.203 0.238 0.194 0.235
192 0.217 0.264 0.235 0.278 0.250 0.278 0.247 0.265 0.271 0.259 0.246 0.251 - - 0.278 0.329 0.263 0.294 0.256 0.290 0.249 0.285
336 0.285 0.312 0.290 0.321 0.304 0.340 0.283 0.303 0.286 0.297 0.274 0.291 - - 0.306 0.346 0.321 0.339 0.314 0.336 0.302 0.327
720 0.359 0.348 0.351 0.383 0.353 0.395 0.373 0.354 0.373 0.354 0.337 0.340 - - 0.350 0.374 0.404 0.397 0.397 0.396 0.372 0.378
Avg. 0.255 0.284 0.262 0.300 0.277 0.305 0.275 0.286 0.287 0.281 0.264 0.273 - - 0.294 0.326 0.300 0.318 0.292 0.315 0.279 0.306
$1^{\text{st}}$ Count
17 6 1 4 9 8 1 3 0 0 3
Discussion on Larger Time Series Models
[…]
$\text{IGC}=\frac{1}{\text{MSE}\times\text{Parameters Count}^{\alpha}\times\text{Training Data Size}^{\beta}},$ (17)
[…]
Table 15: Zero-shot long-term forecasting performance comparison with larger-scale time series forecasting foundation models. A lower MSE or MAE indicates a better prediction. A higher IGC indicates better efficiency. Bold: the best, Underline: the second best.
Models
Comparison with Larger Forecasting Models
FeDaL (Ours) Time-MoEbase Time-MoElarge Time-MoEultra
Metrics
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.347 0.381 0.357 0.381 0.350 0.382 0.349 0.379
192 0.398 0.410 0.384 0.404 0.388 0.412 0.395 0.413
336 0.425 0.452 0.411 0.434 0.411 0.430 0.447 0.453
720 0.457 0.469 0.449 0.477 0.427 0.455 0.457 0.462
Avg. 0.407 0.429 0.400 0.424 0.394 0.419 0.412 0.426
ETTh2 96 0.307 0.355 0.305 0.359 0.302 0.354 0.292 0.352
192 0.349 0.372 0.351 0.386 0.364 0.385 0.347 0.379
336 0.387 0.395 0.391 0.418 0.417 0.425 0.406 0.419
720 0.401 0.406 0.419 0.454 0.537 0.496 0.439 0.447
Avg. 0.361 0.382 0.366 0.404 0.405 0.415 0.371 0.399
ETTm1 96 0.289 0.346 0.338 0.368 0.309 0.557 0.281 0.341
192 0.317 0.369 0.353 0.388 0.346 0.381 0.305 0.358
336 0.370 0.420 0.381 0.413 0.373 0.408 0.369 0.395
720 0.464 0.426 0.504 0.493 0.475 0.477 0.469 0.472
Avg. 0.360 0.390 0.394 0.415 0.376 0.405 0.356 0.391
ETTm2 96 0.207 0.283 0.201 0.291 0.197 0.286 0.198 0.288
192 0.248 0.333 0.258 0.334 0.250 0.322 0.235 0.312
336 0.316 0.340 0.324 0.373 0.337 0.375 0.293 0.348
720 0.397 0.408 0.488 0.464 0.480 0.461 0.427 0.423
Avg. 0.292 0.341 0.317 0.365 0.316 0.361 0.288 0.344
Weather 96 0.159 0.212 0.160 0.214 0.159 0.213 0.157 0.211
192 0.217 0.264 0.210 0.260 0.215 0.266 0.208 0.256
336 0.285 0.312 0.274 0.309 0.291 0.322 0.255 0.290
720 0.359 0.348 0.418 0.405 0.415 0.400 0.405 0.397
Avg. 0.255 0.284 0.265 0.297 0.270 0.300 0.256 0.288
Average
0.335 0.370 0.343 0.382 0.355 0.387 0.342 0.369
$1^{\text{st}}$ Count
21 3 7 22
Total Param.#
28.42 M 113 M 453 M 2.4 B
Training Data
$\sim$ 231B 300B 300B 300B
Information Gain Per Cost
$\mathbf{4.545\times 10^{-19}}$ $8.605\times 10^{-20}$ $2.073\times 10^{-20}$ $4.061\times 10^{-21}$
C.2 Short-term Forecasting
The full short-term forecasting results are presented in Table˜16 and Table˜17. Specifically, Table˜16 compares our FeDaL-pretrained TSFM with advanced deep time series models, while Table˜17 focuses on comparisons among FL-based TSFM pretraining methods. Across all settings, FeDaL consistently outperforms all baselines, including task-specific deep models, general-purpose foundation models, and alternative federated pretraining strategies, highlighting its effectiveness in learning robust and generalizable temporal representations under decentralization.
Table 16: Short-term forecasting results. Bold: the best, Underline: the second best.
Intervals Methods FeDaL (Ours) Time-LLM GPT4TS TimesNet PatchTST N-HiTS N-BEATS ETS.* LightTS DLinear FED.* Stationary Auto.* In.* Re.*
Yearly SMAPE 13.102 13.419 15.11 15.378 13.477 13.422 13.487 18.009 14.247 16.965 14.021 13.717 13.974 14.727 16.169
MASE 2.812 3.005 3.565 3.554 3.019 3.056 3.036 4.487 3.109 4.283 3.036 3.078 3.134 3.418 3.800
OWA 0.748 0.789 0.911 0.918 0.792 0.795 0.795 1.115 0.827 1.058 0.811 0.807 0.822 0.881 0.973
Quarterly SMAPE 9.808 10.110 10.597 10.465 10.38 10.185 10.564 13.376 11.364 12.145 11.1 10.958 11.338 11.360 13.313
MASE 1.112 1.178 1.253 1.227 1.233 1.18 1.252 1.906 1.328 1.520 1.35 1.325 1.365 1.401 1.775
OWA 0.847 0.889 0.938 0.923 0.921 0.893 0.936 1.302 1.000 1.106 0.996 0.981 1.012 1.027 1.252
Monthly SMAPE 12.124 12.980 13.258 13.513 12.959 13.059 13.089 14.588 14.014 13.514 14.403 13.917 13.958 14.062 20.128
MASE 0.898 0.963 1.003 1.039 0.970 1.013 0.996 1.368 1.053 1.037 1.147 1.097 1.103 1.141 2.614
OWA 0.820 0.903 0.931 0.957 0.905 0.929 0.922 1.149 0.981 0.956 1.038 0.998 1.002 1.024 1.927
Others SMAPE 4.508 4.795 6.124 6.913 4.952 4.711 6.599 7.267 15.880 6.709 7.148 6.302 5.485 24.460 32.491
MASE 2.890 3.178 4.116 4.507 3.347 3.054 4.43 5.240 11.434 4.953 4.041 4.064 3.865 20.960 33.355
OWA 0.973 1.006 1.259 1.438 1.049 0.977 1.393 1.591 3.474 1.487 1.389 1.304 1.187 5.879 8.679
Average SMAPE 11.412 11.983 12.69 12.88 12.059 12.035 12.25 14.718 13.525 13.639 13.16 12.780 12.909 14.086 18.200
MASE 1.489 1.595 1.808 1.836 1.623 1.625 1.698 2.408 2.111 2.095 1.775 1.756 1.771 2.718 4.223
OWA 0.818 0.859 0.94 0.955 0.869 0.869 0.896 1.172 1.051 1.051 0.949 0.930 0.939 1.230 1.775
Table 17: Short-term forecasting results based on FL. Bold: the best, Underline: the second best.
Setting Federated Foundation Models Foundation Models
Intervals Methods FeDaL (Ours) FFTS FedAvg MOMENT
Yearly SMAPE 13.102 13.289 14.784 20.649
MASE 2.812 2.909 3.257 4.757
OWA 0.748 0.781 0.866 1.230
Quarterly SMAPE 9.808 10.005 10.920 10.849
MASE 1.112 1.190 1.367 1.305
OWA 0.847 0.877 0.957 0.968
Monthly SMAPE 12.124 11.920 13.048 14.497
MASE 0.898 0.879 1.027 1.143
OWA 0.820 0.815 0.916 1.040
Others SMAPE 4.508 4.490 5.210 5.634
MASE 2.890 2.907 3.657 4.102
OWA 0.973 0.994 1.154 1.240
Average SMAPE 11.412 11.404 12.342 14.593
MASE 1.489 1.522 1.753 2.161
OWA 0.818 0.831 0.926 1.103
Table 18: Full results on imputation task. Bold: the best, Underline: the second best.
Dataset Method FeDaL (Ours) FFTS FedAvg GPT4TS DLinear PatchTST TimesNet FED. Auto. Stationary ETS. LightTS In. Re.
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 12.5% 0.013 0.073 0.015 0.075 0.020 0.105 0.017 0.085 0.023 0.101 0.041 0.130 0.096 0.229 0.093 0.206 0.080 0.193 0.052 0.166 0.032 0.119 0.046 0.144 0.063 0.180 0.042 0.146
25% 0.013 0.071 0.015 0.081 0.022 0.109 0.022 0.096 0.023 0.101 0.044 0.135 0.096 0.229 0.093 0.206 0.080 0.193 0.052 0.166 0.032 0.119 0.046 0.144 0.063 0.180 0.042 0.146
37.5% 0.033 0.091 0.036 0.094 0.046 0.131 0.029 0.111 0.029 0.111 0.049 0.143 0.133 0.271 0.113 0.231 0.103 0.219 0.069 0.191 0.039 0.131 0.057 0.161 0.079 0.200 0.063 0.182
50% 0.030 0.110 0.030 0.111 0.048 0.135 0.040 0.128 0.036 0.124 0.055 0.151 0.186 0.323 0.134 0.255 0.132 0.248 0.089 0.218 0.047 0.145 0.067 0.174 0.093 0.218 0.082 0.208
Avg 0.022 0.090 0.024 0.093 0.034 0.120 0.028 0.105 0.027 0.107 0.047 0.140 0.120 0.253 0.104 0.218 0.093 0.206 0.062 0.177 0.036 0.126 0.051 0.150 0.071 0.188 0.055 0.166
ETTh2 12.5% 0.014 0.062 0.014 0.070 0.020 0.085 0.017 0.076 0.018 0.080 0.026 0.094 0.108 0.239 0.034 0.127 0.062 0.166 0.056 0.159 0.021 0.088 0.023 0.092 0.133 0.270 0.108 0.228
25% 0.017 0.070 0.017 0.074 0.024 0.090 0.020 0.080 0.020 0.085 0.028 0.099 0.164 0.294 0.042 0.143 0.085 0.196 0.080 0.195 0.024 0.096 0.026 0.101 0.135 0.272 0.136 0.262
37.5% 0.018 0.070 0.017 0.078 0.028 0.106 0.022 0.087 0.023 0.091 0.030 0.104 0.237 0.356 0.051 0.159 0.106 0.222 0.110 0.231 0.027 0.103 0.030 0.108 0.155 0.293 0.175 0.300
50% 0.022 0.079 0.019 0.086 0.032 0.111 0.025 0.095 0.026 0.098 0.034 0.110 0.323 0.421 0.059 0.174 0.131 0.247 0.156 0.276 0.030 0.108 0.035 0.119 0.200 0.333 0.211 0.329
Avg 0.018 0.071 0.017 0.074 0.026 0.098 0.021 0.084 0.022 0.088 0.029 0.102 0.208 0.327 0.046 0.151 0.096 0.208 0.101 0.215 0.026 0.099 0.029 0.105 0.156 0.292 0.157 0.280
ETTm1 12.5% 0.030 0.116 0.034 0.132 0.038 0.132 0.043 0.140 0.057 0.159 0.093 0.201 0.126 0.263 0.240 0.345 0.151 0.267 0.070 0.190 0.060 0.165 0.074 0.182 0.114 0.234 0.074 0.194
25% 0.038 0.130 0.041 0.147 0.046 0.142 0.054 0.156 0.069 0.178 0.107 0.217 0.169 0.304 0.265 0.364 0.180 0.292 0.106 0.236 0.080 0.189 0.090 0.203 0.140 0.262 0.102 0.227
37.5% 0.061 0.150 0.065 0.160 0.058 0.162 0.072 0.180 0.084 0.196 0.120 0.230 0.220 0.347 0.296 0.382 0.215 0.318 0.124 0.258 0.102 0.212 0.109 0.222 0.174 0.293 0.135 0.261
50% 0.086 0.192 0.091 0.202 0.074 0.180 0.107 0.216 0.102 0.215 0.141 0.248 0.293 0.402 0.334 0.404 0.257 0.347 0.165 0.299 0.133 0.240 0.137 0.248 0.215 0.325 0.179 0.298
Avg 0.054 0.147 0.058 0.160 0.054 0.154 0.069 0.173 0.078 0.187 0.115 0.224 0.202 0.329 0.284 0.373 0.201 0.306 0.117 0.246 0.094 0.201 0.103 0.214 0.161 0.279 0.122 0.245
ETTm2 12.5% 0.020 0.100 0.028 0.102 0.045 0.128 0.039 0.125 0.040 0.130 0.057 0.152 0.187 0.319 0.101 0.231 0.100 0.216 0.095 0.212 0.042 0.133 0.044 0.138 0.305 0.431 0.163 0.289
25% 0.021 0.104 0.034 0.133 0.052 0.153 0.044 0.135 0.046 0.141 0.061 0.158 0.279 0.390 0.115 0.246 0.127 0.247 0.137 0.258 0.049 0.147 0.050 0.149 0.322 0.444 0.206 0.331
37.5% 0.048 0.110 0.063 0.153 0.077 0.175 0.051 0.147 0.052 0.151 0.067 0.166 0.400 0.465 0.126 0.257 0.158 0.276 0.187 0.304 0.056 0.158 0.060 0.163 0.353 0.462 0.252 0.370
50% 0.048 0.118 0.033 0.151 0.075 0.160 0.059 0.158 0.060 0.162 0.073 0.174 0.602 0.572 0.136 0.268 0.183 0.299 0.232 0.341 0.065 0.170 0.068 0.173 0.369 0.472 0.316 0.419
Avg 0.034 0.108 0.046 0.135 0.062 0.154 0.048 0.141 0.049 0.146 0.065 0.163 0.367 0.436 0.119 0.250 0.142 0.259 0.163 0.279 0.053 0.152 0.055 0.156 0.337 0.452 0.234 0.352
Weather 12.5% 0.013 0.036 0.020 0.053 0.020 0.053 0.026 0.049 0.025 0.045 0.029 0.049 0.057 0.141 0.047 0.101 0.039 0.084 0.041 0.107 0.027 0.051 0.026 0.047 0.037 0.093 0.031 0.076
25% 0.018 0.025 0.024 0.057 0.024 0.057 0.028 0.052 0.029 0.052 0.031 0.053 0.065 0.155 0.052 0.111 0.048 0.103 0.064 0.163 0.029 0.056 0.030 0.054 0.042 0.100 0.035 0.082
37.5% 0.025 0.060 0.034 0.063 0.034 0.063 0.033 0.060 0.031 0.057 0.035 0.058 0.081 0.180 0.058 0.121 0.057 0.117 0.107 0.229 0.033 0.062 0.032 0.060 0.049 0.111 0.040 0.091
50% 0.027 0.050 0.038 0.061 0.038 0.061 0.037 0.065 0.034 0.062 0.038 0.063 0.102 0.207 0.065 0.133 0.066 0.134 0.183 0.312 0.037 0.068 0.037 0.067 0.053 0.114 0.046 0.099
Avg 0.024 0.048 0.029 0.058 0.034 0.050 0.031 0.056 0.030 0.054 0.060 0.144 0.076 0.171 0.055 0.117 0.052 0.110 0.099 0.203 0.032 0.059 0.031 0.057 0.045 0.104 0.038 0.087
$1^{\text{st}}$ Count
40 7 4 1 2 0 0 0 0 0 0 0 0 0
C.3 Time Series Imputation
The full imputation results are presented in Table˜18. Our proposed FeDaL consistently outperforms both advanced deep time series models and federated TSFM pretraining baselines, demonstrating superior generalization across heterogeneous input gaps.
Table 19: Full results of the anomaly detection task. The P, R, and F1 represent the precision, recall, and F1-score (%) respectively. F1-score is the harmonic mean of precision and recall. A higher value of P, R, and F1 indicates a better performance. Bold: the best, Underline: the second best.
Dataset
SMD
MSL
SMAP
SWaT
PSM
Avg. F1
Metric
P
R
F1
P
R
F1
P
R
F1
P
R
F1
P
R
F1
LSTM
[33]
78.52
65.47
71.41
78.04
86.22
81.93
91.06
57.49
70.48
78.06
91.72
84.34
69.24
99.53
81.67
77.97
Transformer
[60]
83.58
76.13
79.56
71.57
87.37
78.68
89.37
57.12
69.70
68.84
96.53
80.37
62.75
96.56
76.07
76.88
LogTrans
[39]
83.46
70.13
76.21
73.05
87.37
79.57
89.15
57.59
69.97
68.67
97.32
80.52
63.06
98.00
76.74
76.60
TCN
[26]
84.06
79.07
81.49
75.11
82.44
78.60
86.90
59.23
70.45
76.59
95.71
85.09
54.59
99.77
70.57
77.24
Reformer
[36]
82.58
69.24
75.32
85.51
83.31
84.40
90.91
57.44
70.40
72.50
96.53
82.80
59.93
95.38
73.61
77.31
Informer
[80]
86.60
77.23
81.65
81.77
86.48
84.06
90.11
57.13
69.92
70.29
96.75
81.43
64.27
96.33
77.10
78.83
Anomaly
[69]
88.91
82.23
85.49
79.61
87.37
83.31
91.85
58.11
71.18
72.51
97.32
83.10
68.35
94.72
79.40
80.50
Pyraformer
[45]
85.61
80.61
83.04
83.81
85.93
84.86
92.54
57.71
71.09
87.92
96.00
91.78
71.67
96.02
82.08
82.57
Autoformer
[68]
88.06
82.35
85.11
77.27
80.92
79.05
90.40
58.62
71.12
89.85
95.81
92.74
99.08
88.15
93.29
84.26
LSSL
[29]
78.51
65.32
71.31
77.55
88.18
82.53
89.43
53.43
66.90
79.05
93.72
85.76
66.02
92.93
77.20
76.74
NSformer
[47]
88.33
81.21
84.62
68.55
89.14
77.50
89.37
59.02
71.09
68.03
96.75
79.88
97.82
96.76
97.29
82.08
DLinear
[73]
83.62
71.52
77.10
84.34
85.42
84.88
92.32
55.41
69.26
80.91
95.30
87.52
98.28
89.26
93.55
82.46
ETSformer
[65]
87.44
79.23
83.13
85.13
84.93
85.03
92.25
55.75
69.50
90.02
80.36
84.91
99.31
85.28
91.76
82.87
LightTS
[78]
87.10
78.42
82.53
82.40
75.78
78.95
92.58
55.27
69.21
91.98
94.72
93.33
98.37
95.97
97.15
84.23
FEDformer
[81]
87.95
82.39
85.08
77.14
80.07
78.57
90.47
58.10
70.76
90.17
96.42
93.19
97.31
97.16
97.23
84.97
TimesNet
[66]
88.66
83.14
85.81
83.92
86.42
85.15
92.52
58.29
71.52
86.76
97.32
91.74
98.19
96.76
97.47
86.34
ModernTCN
[51]
87.86
83.85
85.81
83.94
85.93
84.92
93.17
57.69
71.26
91.83
95.98
93.86
98.09
96.38
97.23
86.62
MOMENT
[28]
78.88
92.01
84.94
88.98
75.10
81.45
90.02
56.51
69.43
92.13
91.67
91.90
98.82
89.55
93.96
84.34
GPT4TS
[82]
88.89
84.98
86.89
82.00
82.91
82.45
90.60
60.95
72.88
92.20
96.34
94.23
98.62
95.68
97.13
86.72
FedAvg
[53]
87.88
89.21
88.44
81.69
83.00
82.32
89.26
58.66
70.78
90.27
90.43
90.32
96.56
95.08
95.86
85.54
FFTS
[9]
89.26
90.48
89.88
89.23
87.64
88.42
90.64
58.88
71.38
91.22
90.99
91.12
99.00
97.85
98.54
87.86
FeDaL
89.15
87.78
88.46
88.87
89.23
89.05
88.46
60.29
71.70
92.56
98.49
95.40
99.29
98.46
98.88
88.70
C.4 Time Series Anomaly Detection
The full anomaly detection results are presented in Table˜19. FeDaL again achieves the best performance among all evaluated methods, surpassing both advanced deep time series models and existing federated TSFM pretraining approaches.
C.5 Federated Scaling Behaviors
While prior work explores scaling laws of centralized TSFMs in terms of model and data size, we instead investigate scaling behaviors in the federated setting, focusing on three key factors: (i) pretraining data size, (ii) number of clients, and (iii) client participation rate. Specifically, we vary: (i) Data size from $\{40\text{B},80\text{B},120\text{B},160\text{B},200\text{B},231\text{B}\}$ with a fixed client count of 174; (ii) Number of clients from $\{30,70,110,140,174\}$ using the full 231B dataset; (iii) Client participation rate from $\{10\%,30\%,50\%,70\%,100\%\}$ with the original pretraining setup. All experiments follow the training protocol in Section 4.3. Results are presented in Figure˜9. We observe that: (1) Increasing data size consistently improves downstream performance, even with a fixed client count; (2) More clients (with constant total data size) lead to better representations, suggesting improved learning of diverse, domain-specific patterns; (3) Higher client participation rates yield stronger results, likely due to more effective aggregation and reduced drift. These findings suggest that federated TSFM pretraining benefits from scaling in data and client dimensions, much like centralized pretraining. However, unlike the “scale-is-all” mindset focused on model size, our results highlight the importance of expanding dataset coverage and client participation as more efficient and federated-aligned strategies for improving generalization.
Figure 9: Scaling behavior from a federated learning perspective. Note that for forecasting and imputation, the y-axis is inverted for better visualization (lower values indicate better performance).